### v2

# Data-driven Implementations of Various Generalizations of Balanced Truncation

Abstract
There exist two main approaches for non-intrusive implementations of approximate balanced truncation within the Loewner framework: the quadrature-based method [1] and the Alternating Direction Implicit (ADI)-based method [2]. Both approaches rely solely on samples of the transfer function to construct truncated balanced models, eliminating the need for access to the original model’s state-space realization. Recently, the quadrature-based approach has been extended to various generalizations of balanced truncation, including positive-real balanced truncation, bounded-real balanced truncation, and balanced stochastic truncation. While this extension [3] is theoretically non-intrusive—meaning it does not require the original state-space realization—it depends on samples of spectral factorizations of the transfer function. Since practical methods for obtaining such samples are currently unavailable, this extension remains largely a theoretical contribution.
In this work, we present a non-intrusive ADI-type framework for these generalized balanced truncation methods that requires only samples of the original transfer function for implementation. To achieve this, we first demonstrate how the approximations of ADI-type method for Riccati equations (RADI) [4] can be equivalently obtained via projection-based interpolation. Leveraging this insight, we develop several non-intrusive, data-driven algorithms that not only satisfy the desired interpolation conditions but also preserve key system properties such as positive-realness, bounded-realness, and minimum phase. Notably, these algorithms rely solely on solving Lyapunov and Sylvester equations, avoiding computationally intensive Riccati equations or linear matrix inequalities. Additionally, we propose non-intrusive, data-driven implementations of RADI (and ADI)-based generalizations of balanced truncation, covering Linear Quadratic Gaussian (LQG) balanced truncation, $\mathcal{H}_{\infty}$ balanced truncation, positive-real balanced truncation, bounded-real balanced truncation, self-weighted balanced truncation, and balanced stochastic truncation. We also discuss practical considerations for improving computational efficiency. Numerical experiments demonstrate that the proposed algorithms perform comparably to their intrusive counterparts.
keywords:
ADI, Balanced truncation, Data-driven, Low-rank, Non-intrusive
journal: ArXiv.org
1 Introduction
Model order reduction (MOR) refers to system-theoretic techniques used to obtain simplified models that retain the essential input-output behavior of high-dimensional dynamical systems. These reduced order models (ROMs) are constructed to preserve the key dynamics and properties of the original system while being much lower in dimension. As a result, they offer significant computational advantages, making simulation, control, and analysis more efficient. For a detailed overview of MOR methods, refer to [5, 6, 7, 8].
Among MOR techniques, balanced truncation (BT) is a widely adopted and effective approach for linear systems [9]. It ensures that the ROM preserves the stability of the original system and provides guaranteed a priori error bound. BT truncates states that contribute minimally to the system’s input-output behavior, as determined by the system’s Hankel singular values. This results in a ROM that closely approximates the full-order model.
The main computational burden in BT lies in solving high-order Lyapunov equations to obtain the system Gramians. Numerous efficient algorithms, as discussed in surveys [10, 11], have been proposed for this purpose. These approaches require access to the system’s full state-space model and are therefore considered intrusive. In contrast, non-intrusive methods do not rely on internal state-space information but instead use system response data, such as transfer function samples or impulse responses [12, 13, 14, 15]. A notable example is the Quadrature-BT (QuadBT) algorithm introduced in [1], which constructs the ROM using numerical integration based on transfer function evaluations along the $j\omega$-axis or on sampled impulse responses and their derivatives. Over the last twenty years, the low-rank Alternating Direction Implicit (ADI) method has become a powerful tool for reducing the computational cost of BT in large-scale settings [16]. Recently, a data-driven, non-intrusive implementation of the ADI-based BT, which constructs the ROM using transfer function samples taken from the right half of the s-plane, was proposed in [2].
BT has been generalized to various classes of dynamical systems [17], including descriptor systems [18], parametric systems [19], nonlinear systems [20], and second-order systems [21], while maintaining important properties such as passivity [22], contractivity [23], and the minimum-phase condition [24]. Recently, QuadBT was extended to include certain generalizations of BT, such as balanced stochastic truncation (BST) [24], positive-real BT (PR-BT) [24], and bounded-real BT (BR-BT) [23, 25]. Although this extension [3] is theoretically non-intrusive, it relies on access to samples of spectral factorizations of the transfer function. As practical techniques for acquiring such samples are not yet available, the method remains primarily of theoretical interest.
In this work, we introduce non-intrusive, data-driven interpolation algorithms that not only interpolate the original transfer function at chosen interpolation points but also preserve key system properties such as positive-realness, bounded-realness, and the minimum-phase condition. Additionally, we present an ADI-based non-intrusive data-driven framework applicable to several BT generalizations, including Linear Quadratic Gaussian-BT (LQG-BT) [26], $\mathcal{H}_{\infty}$-BT [27], PR-BT [24], BR-BT [23], self-weighted BT (SW-BT) [28], and BST [24]. In contrast to the quadrature-based approach in [3], the proposed methods rely on transfer function samples that can be practically obtained using frequency-domain techniques from the system identification literature [29]. The data-driven algorithms for LQG-BT and $\mathcal{H}_{\infty}$-BT proposed in this paper provide a control design approach that can be used to design LQG and $\mathcal{H}_{\infty}$-infinity controllers for unknown plants using only samples of the plant’s transfer function. Numerical results using a benchmark test model demonstrate that the proposed algorithms perform comparably to their intrusive counterparts.
The paper is structured as follows. Section 2 introduces the MOR problem and provides a brief overview of key MOR techniques relevant to this work. Section 3 presents the main contributions. In Subsection 3.1, we propose non-intrusive data-driven implementations of LQG-BT and $\mathcal{H}_{\infty}$-BT. Subsection 3.2 introduces a data-driven implementation of PR-BT. Subsection 3.3 focuses on a data-driven BR-BT implementation. In Subsection 3.4, we propose a data-driven implementation of SW-BT. Subsection 3.5 presents a data-driven implementation of BST. Section 4 addresses practical considerations to prevent numerical issues during implementation. Section 5 evaluates the performance of the proposed algorithms through a benchmark numerical example. Finally, Section 6 offers concluding remarks.
2 Preliminaries
Consider a stable $n^{th}$-order linear time-invariant (LTI) system $G(s)$, represented in minimal state-space form as
$G(s)=C(sI-A)^{-1}B+D,$
with matrices $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, $C\in\mathbb{R}^{p\times n}$, and $D\in\mathbb{R}^{p\times m}$.
Consider an $r^{th}$-order stable reduced-order approximation $\hat{G}(s)$ of $G(s)$ with minimal state-space realization:
$\hat{G}(s)=\hat{C}(sI-\hat{A})^{-1}\hat{B}+D,$
where $\hat{A}\in\mathbb{C}^{r\times r}$, $\hat{B}\in\mathbb{C}^{r\times m}$, $\hat{C}\in\mathbb{C}^{p\times r}$, and $r\ll n$.
The ROM $\hat{G}(s)$ is constructed from $G(s)$ through Petrov-Galerkin projection:
$\hat{A}=\hat{W}^{*}A\hat{V},\quad\hat{B}=\hat{W}^{*}B,\quad\hat{C}=C\hat{V}$,
[…]
2.1 Interpolation Theory [31]
Consider the right interpolation points $(\sigma_{1},\dots,\sigma_{v})$ and left interpolation points $(\mu_{1},\dots,\mu_{w})$. In the interpolation framework, the projection matrices $\hat{V}\in\mathbb{C}^{n\times vm}$ and $\hat{W}\in\mathbb{C}^{n\times wp}$ are constructed as:
$\hat{V}$ $=\begin{bmatrix}(\sigma_{1}I-A)^{-1}B&\cdots&(\sigma_{v}I-A)^{-1}B\end{bmatrix}$, (1)
$\hat{W}$ $=\begin{bmatrix}(\mu_{1}^{*}I-A^{T})^{-1}C^{T}&\cdots&(\mu_{w}^{*}I-A^{T})^{-1}C^{T}\end{bmatrix}$. (2)
To enforce the Petrov-Galerkin condition ($\hat{W}^{*}\hat{V}=I$), $\hat{W}$ is replaced with $\hat{W}T_{w}$, where $T_{w}=(\hat{V}^{*}\hat{W})^{-1}$. The resulting ROM satisfies the interpolation conditions:
$G(\sigma_{i})=G_{r}(\sigma_{i}),\quad G(\mu_{i})=G_{r}(\mu_{i}).$ (3)
Additionally, if there are common right and left interpolation points, i.e., $\sigma_{i}=\mu_{j}$, the following Hermite interpolation conditions are also satisfied for those points:
$G^{\prime}(\sigma_{i})=G_{r}^{\prime}(\sigma_{i}).$ (4)
2.2 Interpolatory Loewner framework [12]
The matrix $D$ can be obtained by sampling the transfer function $G(s)$ at a sufficiently high frequency, since $D=\lim_{s\to\infty}G(s)$. Let $H(s)$ be defined as:
$H(s)=C(sI-A)^{-1}B.$
Samples of $H(s_{i})$ can be computed from those of $G(s_{i})$ as:
$H(s_{i})=G(s_{i})-G(\infty).$
In the Loewner framework, the following matrices are constructed from $H(s)$ samples at interpolation points:
$\hat{W}^{*}\hat{V}$ $=\begin{bmatrix}-\frac{H(\sigma_{1})-H(\mu_{1})}{\sigma_{1}-\mu_{1}}&\cdots&-\frac{H(\sigma_{v})-H(\mu_{1})}{\sigma_{v}-\mu_{1}}\\ \vdots&\ddots&\vdots\\ -\frac{H(\sigma_{1})-H(\mu_{w})}{\sigma_{1}-\mu_{w}}&\cdots&-\frac{H(\sigma_{v})-H(\mu_{w})}{\sigma_{v}-\mu_{w}}\end{bmatrix}$,
$\hat{W}^{*}A\hat{V}$ $=\begin{bmatrix}-\frac{\sigma_{1}H(\sigma_{1})-\mu_{1}H(\mu_{1})}{\sigma_{1}-\mu_{1}}&\cdots&-\frac{\sigma_{v}H(\sigma_{v})-\mu_{1}H(\mu_{1})}{\sigma_{v}-\mu_{1}}\\ \vdots&\ddots&\vdots\\ -\frac{\sigma_{1}H(\sigma_{1})-\mu_{w}H(\mu_{w})}{\sigma_{1}-\mu_{w}}&\cdots&-\frac{\sigma_{v}H(\sigma_{v})-\mu_{w}H(\mu_{w})}{\sigma_{v}-\mu_{w}}\end{bmatrix}$,
$\hat{W}^{*}B$ $=\begin{bmatrix}H(\mu_{1})\\ \vdots\\ H(\mu_{w})\end{bmatrix},\quad C\hat{V}=\begin{bmatrix}H(\sigma_{1})&\cdots H(\sigma_{v})\end{bmatrix}$, (5)
where $\hat{V}$ and $\hat{W}$ are as defined in (1) and (2), respectively. When $\sigma_{j}\approx\mu_{i}$,
$\frac{H(\sigma_{j})-H(\mu_{i})}{\sigma_{j}-\mu_{i}}$ $\approx H^{\prime}(\sigma_{j}),$
$\frac{\sigma_{j}H(\sigma_{j})-\mu_{i}H(\mu_{i})}{\sigma_{j}-\mu_{i}}$ $\approx H(\sigma_{j})+\sigma_{j}H^{\prime}(\sigma_{j}).$
Thus, if the sets of right and left interpolation points share common elements, derivative samples $H^{\prime}(\sigma_{j})$ are also needed to construct $\hat{W}^{*}\hat{V}$ and $\hat{W}^{*}A\hat{V}$. The matrices $\hat{W}^{*}\hat{V}$ and $\hat{W}^{*}A\hat{V}$ exhibit a special structure known as the Loewner matrix and shifted Loewner matrix, respectively, which gives the name “Interpolatory Loewner Framework”. The matrix $\hat{W}^{*}\hat{V}$ is not always invertible, preventing the computation of $\hat{A}=(\hat{W}^{*}\hat{V})^{-1}\hat{W}^{*}A\hat{V}$ and $\hat{B}=(\hat{W}^{*}\hat{V})^{-1}\hat{W}^{*}B$. Instead, the ROM is constructed in descriptor form:
$\hat{G}(s)=C\hat{V}(s\hat{W}^{*}\hat{V}-\hat{W}^{*}A\hat{V})^{-1}\hat{W}^{*}B+G(\infty).$
2.3 Pseudo-optimal Rational Krylov Algorithm (PORK) [32]
The matrices $S_{v}$, $S_{w}$, $L_{v}$, and $L_{w}$ are defined as:
$S_{v}$ $=\text{diag}(\sigma_{1},\dots,\sigma_{v})\otimes I_{m},$ $S_{w}$ $=\text{diag}(\mu_{1},\dots,\mu_{w})\otimes I_{p},$
$L_{v}$ $=\begin{bmatrix}1,\dots,1\end{bmatrix}\otimes I_{m},$ $L_{w}^{T}$ $=\begin{bmatrix}1,\dots,1\end{bmatrix}\otimes I_{p}.$ (6)
The projection matrices $\hat{V}$ and $\hat{W}$ in (1) and (2) solve:
$A\hat{V}-\hat{V}S_{v}+BL_{v}$ $=0,$ (7)
$A^{T}\hat{W}-\hat{W}S_{w}^{*}+C^{T}L_{w}^{T}$ $=0.$ (8)
Pre-multiplying (7) with $\hat{W}^{*}$ under the condition $\hat{W}^{*}\hat{V}=I$ reveals that $\hat{A}=S_{v}-\hat{B}L_{v}$. This parameterizes $A_{r}$ in terms of $\hat{B}=\zeta$ while preserving the interpolation conditions induced by $\hat{V}$, i.e., varying $\zeta$ is equivalent to varying $\hat{W}$.
Let us assume that all interpolation points $(\sigma_{1},\cdots,\sigma_{v})$ lie in the right half of the $s$-plane. Consider the observable pair $(S_{v},L_{v})$ that satisfies the Lyapunov equation:
$-S_{v}^{*}Q_{v}-Q_{v}S_{v}+L_{v}^{T}L_{v}=0.$ (9)
With the free parameter $\zeta=Q_{v}^{-1}L_{v}^{T}$, we obtain $\hat{A}=S_{v}-\hat{B}L_{v}=-Q_{v}^{-1}S_{v}^{*}Q_{v}$. The resulting ROM
$\hat{A}_{r}=-Q_{v}^{-1}S_{v}^{*}Q_{v},\quad\hat{B}=Q_{v}^{-1}L_{v}^{T},\quad\hat{C}=C\hat{V}$,
[…]
Similarly, pre-multiplying (8) with $\hat{V}^{*}$ shows that $\hat{A}$ can also be expressed as $\hat{A}=S_{w}-L_{w}\hat{C}$. This parameterizes $\hat{A}$ in terms of $\hat{C}=\zeta$ while preserving the interpolation conditions induced by $\hat{W}$, since changing $\zeta$ is equivalent to varying $\hat{V}$.
Let us assume that all interpolation points $(\mu_{1},\cdots,\mu_{w})$ lie in the right half of the $s$-plane. For the controllable pair $(S_{w},L_{w})$ that satisfies the Lyapunov equation:
$-S_{w}P_{w}-P_{w}S_{w}^{*}+L_{w}L_{w}^{T}=0,$ (10)
choosing the free parameter $\zeta=L_{w}^{T}P_{w}^{-1}$ leads to the expression $\hat{A}=S_{w}-L_{w}\hat{C}=-P_{w}S_{w}^{*}P_{w}^{-1}$. The resulting ROM
$\hat{A}=-P_{w}S_{w}^{*}P_{w}^{-1},\quad\hat{B}=\hat{W}^{*}B,\quad\hat{C}=L_{w}^{T}P_{w}^{-1},$ (11)
satisfies the $\mathcal{H}_{2}$-optimality condition given by $\frac{\partial}{\partial\hat{B}}\big(\|G(s)-\hat{G}(s)\|_{\mathcal{H}_{2}}^{2}\big)=0$. This approach is referred to as Output PORK (O-PORK) throughout this paper.
2.4 Balanced Truncation (BT) Family
The controllability Gramian $P$ and observability Gramian $Q$ are defined through the frequency-domain integrals:
$P$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}(j\omega I-A)^{-1}BB^{T}(-j\omega I-A^{T})^{-1}\,d\omega,$ (12)
$Q$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}(-j\omega I-A^{T})^{-1}C^{T}C(j\omega I-A)^{-1}\,d\omega.$ (13)
These Gramians solve the following Lyapunov equations:
$AP+PA^{T}+BB^{T}=0,$ (14)
$A^{T}Q+QA+C^{T}C=0.$ (15)
The Cholesky factorizations of the Gramians are given by
$P=L_{p}L_{p}^{T}\quad\text{and}\quad Q=L_{q}L_{q}^{T}.$
The balanced square-root algorithm (BSA) [33] is among the most numerically stable methods for implementing BT [9]. It begins by computing the singular value decomposition (SVD) of the product $L_{q}^{T}L_{p}$:
$L_{q}^{T}L_{p}=\begin{bmatrix}U_{1}&U_{2}\end{bmatrix}\begin{bmatrix}\Sigma_{r}&0\\ 0&\Sigma_{n-r}\end{bmatrix}\begin{bmatrix}V_{1}^{T}\\ V_{2}^{T}\end{bmatrix}$.
The projection matrices $\hat{W}$ and $\hat{V}$ in BSA are constructed as:
$\hat{W}=L_{q}U_{1}\Sigma_{r}^{-\frac{1}{2}}\quad\text{and}\quad\hat{V}=L_{p}V_{1}\Sigma_{r}^{-\frac{1}{2}}.$
The ROM constructed by BT preserves the stability of the original system $G(s)$ as well as its $r$ largest Hankel singular values $\sqrt{\lambda_{i}(PQ)}$. To preserve additional properties—such as stable minimum phase, positive-realness, or bounded-realness—generalized versions of BT have been developed, which primarily differ in their definitions of the Gramians.
In LQG-BT [26], the Gramian-like matrices are computed by solving the following filter and controller Riccati equations:
$AP_{\text{LQG}}+P_{\text{LQG}}A^{T}+BB^{T}-P_{\text{LQG}}C^{T}CP_{\text{LQG}}=0,$ (16)
$A^{T}Q_{\text{LQG}}+Q_{\text{LQG}}A+C^{T}C-Q_{\text{LQG}}BB^{T}Q_{\text{LQG}}=0.$ (17)
By replacing $P$ and $Q$ with $P_{\text{LQG}}$ and $Q_{\text{LQG}}$ in BT, respectively, a ROM can be obtained that is suitable for designing a reduced-order LQG controller with good closed-loop performance when used with the original plant.
In $\mathcal{H}_{\infty}$-BT [27], the Gramian-like matrices are computed by solving the following filter and controller Riccati equations with $\gamma>0$:
$AP_{\mathcal{H}_{\infty}}+P_{\mathcal{H}_{\infty}}A^{T}+BB^{T}-(1-\gamma^{-2})P_{\mathcal{H}_{\infty}}C^{T}CP_{\mathcal{H}_{\infty}}=0,$ (18)
$A^{T}Q_{\mathcal{H}_{\infty}}+Q_{\mathcal{H}_{\infty}}A+C^{T}C-(1-\gamma^{-2})Q_{\mathcal{H}_{\infty}}BB^{T}Q_{\mathcal{H}_{\infty}}=0.$ (19)
By replacing $P$ and $Q$ with $P_{\mathcal{H}_{\infty}}$ and $Q_{\mathcal{H}_{\infty}}$ in BT, respectively, a ROM can be obtained that is suitable for designing a reduced-order $\mathcal{H}_{\infty}$ controller with good closed-loop performance when used with the original plant.
In PR-BT [24, 34], the Gramians solve the following Riccati equations:
$AP_{\text{PR}}+P_{\text{PR}}A^{T}+(B-P_{\text{PR}}C^{T})(D+D^{T})^{-1}(B-P_{\text{PR}}C^{T})^{T}=0,$ (20)
$A^{T}Q_{\text{PR}}+Q_{\text{PR}}A+(C-B^{T}Q_{\text{PR}})^{T}(D+D^{T})^{-1}(C-B^{T}Q_{\text{PR}})=0.$ (21)
By replacing $P$ and $Q$ with $P_{\text{PR}}$ and $Q_{\text{PR}}$ in BT, respectively, a ROM that preserves the positive-realness of the original model can be obtained.
In BR-BT [35, 34], the Gramians solve the following Riccati equations:
$AP_{\text{BR}}+P_{\text{BR}}A^{T}+BB^{T}+(P_{\text{BR}}C^{T}$ $+BD^{T})(I-DD^{T})^{-1}$
$\times(P_{\text{BR}}C^{T}+BD^{T})^{T}=0,$ (22)
$A^{T}Q_{\text{BR}}+Q_{\text{BR}}A+C^{T}C+(B^{T}Q_{\text{BR}}$ $+D^{T}C)^{T}(I-D^{T}D)^{-1}$
$\times(B^{T}Q_{\text{BR}}+D^{T}C)=0.$ (23)
By replacing $P$ and $Q$ with $P_{\text{BR}}$ and $Q_{\text{BR}}$ in BT, respectively, a ROM that preserves the bounded-realness of the original model can be obtained.
In SW-BT [28], the controllability Gramian $P$ remains the same as in BT, while the weighted observability Gramian $Q_{\text{SW}}$ solves the following Lyapunov equation:
$(A-BD^{-1}C)^{T}Q_{\text{SW}}+Q_{\text{SW}}(A-BD^{-1}C)+C^{T}(DD^{T})^{-1}C=0.$ (24)
By replacing $Q$ with $Q_{\text{SW}}$ in BT, a ROM that preserves the minimum-phase property of the original model can be obtained.
In BST [24, 36], the controllability Gramian $P$ remains the same as in BT, while the weighted observability Gramian $Q_{\text{S}}$ solves the following Riccati equation:
$A^{T}Q_{\text{S}}+Q_{\text{S}}A+\big(C-(CP+DB^{T})Q_{\text{S}}\big)^{T}$ $(DD^{T})^{-1}$
$\times\big(C-(CP+DB^{T})Q_{\text{S}}\big)=0.$ (25)
By replacing $Q$ with $Q_{\text{S}}$ in BT, a ROM that preserves the minimum-phase property of the original model can be obtained. Unlike SW-BT, BST can handle non-minimum-phase models as well. Both SW-BT and BST minimize the relative error $G^{-1}(s)(G(s)-\hat{G}(s))$ and tend to ensure uniform accuracy across the entire frequency spectrum.
2.5 Quadrature-based Balanced Truncation (QuadBT) [1]
The integrals (12) and (13) can be approximated numerically using quadrature rules as:
$P$ $\approx\tilde{P}=\sum_{i=1}^{v}w_{p,i}^{2}(j\sigma_{i}I-A)^{-1}BB^{T}(-j\sigma_{i}I-A^{T})^{-1},$
$Q$ $\approx\tilde{Q}=\sum_{i=1}^{w}w_{q,i}^{2}(-j\mu_{i}I-A^{T})^{-1}C^{T}C(j\mu_{i}I-A)^{-1},$
where $\sigma_{i}$ and $\mu_{i}$ are the quadrature nodes, and $w_{p,i}^{2}$ and $w_{q,i}^{2}$ are their corresponding weights. The factorizations $\tilde{P}=\tilde{L}_{p}\tilde{L}_{p}^{*}$ and $\tilde{Q}=\tilde{L}_{q}\tilde{L}_{q}^{*}$ can be expressed as: $\tilde{L}_{p}=\hat{V}\hat{L}_{p}$ and $\tilde{L}_{q}=\hat{W}\hat{L}_{q}$, with $\hat{V}$ and $\hat{W}$ as in (1) and (2), respectively, and
$\hat{L}_{p}$ $=\text{diag}(w_{p,1},\dots,w_{p,v})\otimes I_{m},$
$\hat{L}_{q}$ $=\text{diag}(w_{q,1},\dots,w_{q,w})\otimes I_{p}.$
Here, $\hat{L}_{p}$ and $\hat{L}_{q}$ depend only on the quadrature weights. Moreover, the terms $\hat{W}^{*}\hat{V}$, $\hat{W}^{*}A\hat{V}$, $\hat{W}^{*}B$, $C\hat{V}$, and $D$ can be constructed non-intrusively using samples of $G(s)$ via the Loewner framework. In the BSA, $\tilde{L}_{p}$ and $\tilde{L}_{q}$ replace $L_{p}$ and $L_{q}$, leading to the decomposition:
$\hat{L}_{q}^{*}(\hat{W}^{*}\hat{V})\hat{L}_{p}=\begin{bmatrix}\tilde{U}_{1}&\tilde{U}_{2}\end{bmatrix}\begin{bmatrix}\tilde{\Sigma}_{r}&0\\ 0&\tilde{\Sigma}_{n-r}\end{bmatrix}\begin{bmatrix}\tilde{V}_{1}^{*}\\ \tilde{V}_{2}^{*}\end{bmatrix}$. (26)
The projection matrices $W_{r}$ and $V_{r}$ are then defined as:
$W_{r}=\hat{L}_{q}\tilde{U}_{1}\tilde{\Sigma}_{r}^{-1/2}\quad\text{and}\quad V_{r}=\hat{L}_{p}\tilde{V}_{1}\tilde{\Sigma}_{r}^{-1/2}.$ (27)
Finally, the ROM in QuadBT is constructed as:
$\hat{A}$ $=W_{r}^{*}(\hat{W}^{*}A\hat{V})V_{r},$ $\hat{B}$ $=W_{r}^{*}(\hat{W}^{*}B),$ $\hat{C}$ $=(C\hat{V})V_{r}.$ (28)
In [3], the following observations were made about PRBT, BRBT, and BST:
  1. 1.
    In PR-BT, the Gramians $P_{\text{PR}}$ and $Q_{\text{PR}}$ are respectively the controllability Gramian and observability Gramian associated with state-space realizations of the spectral factorizations of $G(s)+G^{*}(s)$.
  2. 2.
    In BR-BT, the Gramians $P_{\text{BR}}$ and $Q_{\text{BR}}$ are respectively the controllability Gramian and observability Gramian associated with state-space realizations of the spectral factorizations of $I_{m}-G^{*}(s)G(s)$ and $I_{p}-G(s)G^{*}(s)$.
  3. 3.
    In BST, the observability Gramian $Q_{\text{S}}$ is the frequency-weighted observability Gramian with weight associated with the spectral factorization of $G(s)G^{*}(s)$.
  4. 4.
    The QuadBT framework could be extended to PR-BT, BR-BT, and BST if samples were available for the spectral factorizations of $G(s)+G^{*}(s)$, $I_{m}-G^{*}(s)G(s)$, $I_{p}-G(s)G^{*}(s)$, and $G(s)G^{*}(s)$.
  5. 5.
    Practical methods to acquire these samples are not yet available and are left for future work.
We do not discuss these generalizations in detail here because we focus on methods requiring only samples of $G(s)$, for which practical measurement methods exist, unlike for the spectral factorizations of $G(s)+G^{*}(s)$, $I_{m}-G^{*}(s)G(s)$, $I_{p}-G(s)G^{*}(s)$, and $G(s)G^{*}(s)$.
2.6 Data-driven ADI-based Balanced Truncation (DD-ADI-BT) [2]
In [37], it is shown that PORK and the ADI method [16] produce identical approximations of Lyapunov equations when the ADI shifts are selected as the mirror images of the interpolation points. Consequently, the ADI-based approximation of $P$ for the shifts $(-\sigma_{1},\dots,-\sigma_{v})$ can be obtained via I-PORK as $P\approx\tilde{P}=\hat{V}Q_{v}^{-1}\hat{V}^{*}$. Similarly, the approximation of $Q$ for the shifts $(-\mu_{1},\dots,-\mu_{w})$ can be computed via O-PORK as $Q\approx\tilde{Q}=\hat{W}P_{w}^{-1}\hat{W}^{*}$. By factorizing $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$, and defining $\tilde{L}_{p}=\hat{V}\hat{L}_{p}$ and $\tilde{L}_{q}=\hat{W}\hat{L}_{q}$, the approximations $\tilde{P}$ and $\tilde{Q}$ can be expressed in factorized form as $\tilde{P}=\tilde{L}_{p}\tilde{L}_{p}^{*}$ and $\tilde{Q}=\tilde{L}_{q}\tilde{L}_{q}^{*}$. As noted in [2], replacing $L_{p}$ and $L_{q}$ with $\tilde{L}_{p}$ and $\tilde{L}_{q}$ in BSA yields a non-intrusive implementation of BT. This is because $\hat{L}_{p}$ and $\hat{L}_{q}$ depend only on the interpolation points, and the quantities $\hat{W}^{*}\hat{V}$, $\hat{W}^{*}A\hat{V}$, $\hat{W}^{*}B$, $C\hat{V}$, and $D$ can be computed non-intrusively within the Loewner framework. Thus, the ADI-based BT algorithm can be implemented non-intrusively using transfer function samples in the right half-plane, unlike QuadBT, which requires samples along the imaginary axis.
3 Main Work
In this section, we present several new results on data-driven MOR and controller design, outlined as follows:
  1. 1.
    Recall that the interpolant $\hat{G}(s)=C\hat{V}(sI-S_{v}+\zeta L_{v})^{-1}\zeta+D$ interpolates $G(s)$ at $(\sigma_{1},\cdots,\sigma_{v})$, where $\hat{V}$ is defined in (1) and $\zeta$ is a free parameter. We identify a specific choice of $\zeta$ such that the RADI [4]-based approximation of $P_{\text{LQG}}$ for ADI shifts $(-\sigma_{1},\ldots,-\sigma_{v})$ can be produced by solving a Lyapunov equation involving only the interpolation points $(\sigma_{1},\ldots,\sigma_{v})$ and the product $C\hat{V}$.
  2. 2.
    Similarly, for the interpolant $\hat{G}(s)=\zeta(sI-S_{w}+L_{w}\zeta)^{-1}\hat{W}^{*}B+D$, which interpolates $G(s)$ at $(\mu_{1},\ldots,\mu_{w})$ with $\hat{W}$ defined in (2), we provide a specific $\zeta$ such that the RADI-based approximation of $Q_{\text{LQG}}$ for ADI shifts $(-\mu_{1},\ldots,-\mu_{w})$ can be computed by solving a Lyapunov equation involving only the interpolation data $(\mu_{1},\ldots,\mu_{w})$ and $\hat{W}^{*}B$.
  3. 3.
    Using these approximations of $P_{\text{LQG}}$ and $Q_{\text{LQG}}$, we develop a non-intrusive implementation of LQG-BT, which enables LQG controller design based solely on samples of the plant’s transfer function $G(s)$ in the right-half of the $s$-plane.
  4. 4.
    A similar approach is proposed for the non-intrusive implementation of $\mathcal{H}_{\infty}$-BT, which enables $\mathcal{H}_{\infty}$ controller design based solely on samples of the plant’s transfer function $G(s)$ in the right-half of the $s$-plane.
  5. 5.
    The work in [38] presents linear matrix inequalities (LMIs)- and Riccati-based methods for choosing $\zeta$ in $\hat{G}(s)=C\hat{V}(sI-S_{v}+\zeta L_{v})^{-1}\zeta+D$ that preserve positive-realness, bounded-realness, and minimum-phase properties. However, the formulas in [38] are intrusive since they require access to the state-space matrices $(A,B,C,D)$, and they are also computationally expensive. In contrast, we propose non-intrusive methods to compute $\zeta$ in both $\hat{G}(s)=C\hat{V}(sI-S_{v}+\zeta L_{v})^{-1}\zeta+D$ and $\hat{G}(s)=\zeta(sI-S_{w}+L_{w}\zeta)^{-1}\hat{W}^{*}B+D$ that preserve these system properties. These new formulas require only the transfer function samples and can be derived using Lyapunov and Sylvester equations (or even directly from $G(s)$, as discussed later).
  6. 6.
    We also provide non-intrusive formulas for computing $\zeta$ in the above interpolants that yield the same approximations of $P_{\text{PR}}$, $Q_{\text{PR}}$, $P_{\text{BR}}$, $Q_{\text{BR}}$, $Q_{\text{S}}$ as those produced by RADI. Together with the approximation of $Q_{\text{SW}}$ obtained via O-PORK, we construct non-intrusive implementations of PR-BT, BR-BT, SW-BT, and BST that rely solely on transfer function samples in the right-half of the $s$-plane.
3.1 RADI-based Non-intrusive Implementation of LQG-BT
As demonstrated in [37], the ADI method for computing the low-rank solution of $P$ with shifts $(-\sigma_{1},\dots,-\sigma_{v})$ corresponds to a Petrov-Galerkin approach. This method implicitly interpolates at $(\sigma_{1},\dots,\sigma_{v})$ using the projection matrix $\hat{V}$, while the projection matrix $\hat{W}$ is implicitly chosen to satisfy $\hat{W}^{*}\hat{V}=I$, ensuring that the poles of $\hat{A}$ are placed at $(-\sigma_{1}^{*},\dots,-\sigma_{v}^{*})$. Similarly, [39] shows that the low-rank solution of $P_{\text{LQG}}$ obtained via RADI with shifts $(-\sigma_{1},\dots,-\sigma_{v})$ can be interpreted as a Petrov-Galerkin method. Here, interpolation is implicitly performed at $(\sigma_{1},\dots,\sigma_{v})$ using $\hat{V}$, and $\hat{W}$ is selected such that $\hat{W}^{*}\hat{V}=I$, placing the poles of $\hat{A}-\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C}$ at $(-\sigma_{1}^{*},\dots,-\sigma_{v}^{*})$. The matrix $\hat{P}_{\text{LQG}}$ solves the projected Riccati equation:
$\hat{A}\hat{P}_{\text{LQG}}+\hat{P}_{\text{LQG}}\hat{A}^{*}+\hat{B}\hat{B}^{*}-\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C}\hat{P}_{\text{LQG}}=0.$ (29)
In the following theorem, we specify a choice of $\zeta$ for the interpolant $\hat{G}(s)=C\hat{V}(sI-S_{v}+\zeta L_{v})^{-1}\zeta+D$ such that it places the poles of $\hat{A}-\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C}$ at $(-\sigma_{1}^{*},\dots,-\sigma_{v}^{*})$, thereby yielding the same approximation as RADI.
Theorem 3.1.
Let $\hat{V}$ be as defined in (1), with all interpolation points $(\sigma_{1},\dots,\sigma_{v})$ located in the right half of the $s$-plane. Assume further that the pair $(S_{v},L_{v})$ is observable and that $Q_{v}>0$ uniquely solves the Lyapunov equation:
$-S_{v}^{*}Q_{v}-Q_{v}S_{v}+L_{v}^{T}L_{v}+\hat{C}^{*}\hat{C}=0.$ (30)
Then, the ROM $\hat{H}(s)$, defined by
$\hat{A}=S_{v}-\hat{B}L_{v},\quad\hat{B}=Q_{v}^{-1}L_{v}^{T},\quad\hat{C}=C\hat{V}$,
which interpolates $H(s)$ at $(\sigma_{1},\dots,\sigma_{v})$, satisfies the following properties:
  1. 1.
    The matrix $\hat{A}$ equals $Q_{v}^{-1}(-S_{v}^{*}+\hat{C}^{*}\hat{C}Q_{v}^{-1})Q_{v}$.
  2. 2.
    The solution $\hat{P}_{\text{LQG}}$ to the projected Riccati equation (29) is $Q_{v}^{-1}$.
  3. 3.
    The matrix $\hat{A}-\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C}$ is Hurwitz, with eigenvalues at $(-\sigma_{1}^{*},\dots,-\sigma_{v}^{*})$.
Proof.
  1. 1.
    Pre-multiplying (30) by $Q_{v}^{-1}$ yields:
    $-Q_{v}^{-1}S_{v}^{*}Q_{v}-S_{v}$ $+Q_{v}^{-1}L_{v}^{T}L_{v}+Q_{v}^{-1}\hat{C}^{*}\hat{C}=0$
    $S_{v}-\hat{B}L_{v}$ $=-Q_{v}^{-1}S_{v}^{*}Q_{v}+Q_{v}^{-1}\hat{C}^{*}\hat{C}$
    $\hat{A}$ $=Q_{v}^{-1}(-S_{v}^{*}+\hat{C}^{*}\hat{C}Q_{v}^{-1})Q_{v}.$
  2. 2.
    Consider the expression:
    $\hat{A}Q_{v}^{-1}+Q_{v}^{-1}\hat{A}^{*}+\hat{B}\hat{B}^{*}-Q_{v}^{-1}\hat{C}^{*}\hat{C}Q_{v}^{-1}$
    $=-Q_{v}^{-1}S_{v}^{*}+Q_{v}^{-1}\hat{C}^{*}\hat{C}Q_{v}^{-1}-S_{v}Q_{v}^{-1}+Q_{v}^{-1}\hat{C}^{*}\hat{C}Q_{v}^{-1}$
    $\hskip 119.50148pt+Q_{v}^{-1}L_{v}^{T}L_{v}Q_{v}^{-1}-Q_{v}^{-1}\hat{C}^{T}\hat{C}Q_{v}^{-1}$
    $=-Q_{v}^{-1}S_{v}^{*}-S_{v}Q_{v}^{-1}+Q_{v}^{-1}L_{v}^{T}L_{v}Q_{v}^{-1}+Q_{v}^{-1}\hat{C}^{T}\hat{C}Q_{v}^{-1}$
    $=Q_{v}^{-1}\big(-S_{v}^{*}Q_{v}-Q_{v}S_{v}+L_{v}^{T}L_{v}+\hat{C}^{*}\hat{C}\big)Q_{v}^{-1}$
    $=0.$
    This shows that $Q_{v}^{-1}$ satisfies the projected Riccati equation (29).
  3. 3.
    Given $\hat{A}=\hat{P}_{\text{LQG}}(-S_{v}^{*}+\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C})\hat{P}_{\text{LQG}}^{-1}$, the eigenvalues of $\hat{A}$ are equal to those of $S_{v}^{*}+\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C}$. As $S_{v}^{*}$ is Hurwitz, the matrix $\hat{A}-\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C}$ is Hurwitz with eigenvalues at $(-\sigma_{1}^{*},\cdots,-\sigma_{v}^{*})$.
It is clear from Theorem 3.1 that the RADI-based approximation of $P_{\text{LQG}}$ is given by $P_{\text{LQG}}\approx\hat{V}Q_{v}^{-1}\hat{V}^{*}$.
Note that $Q_{\text{LQG}}$ is the dual of $P_{\text{LQG}}$. For completeness, we present the dual of Theorem 3.1 below, which produces the same approximation of $Q_{\text{LQG}}$ as RADI when the shifts are chosen as $(-\mu_{1},\cdots,-\mu_{w})$.
Theorem 3.2.
Let $\hat{W}$ be as defined in (2), with all interpolation points $(\mu_{1},\dots,\mu_{w})$ located in the right half of the $s$-plane. Assume further that the pair $(S_{w},L_{w})$ is controllable and that $P_{w}>0$ uniquely solves the Lyapunov equation:
$-S_{w}P_{w}-P_{w}S_{w}^{*}+L_{w}L_{w}^{T}+\hat{B}\hat{B}^{*}=0.$ (31)
Then, the ROM $\hat{H}(s)$, defined by
$\hat{A}=S_{w}-L_{w}\hat{C},\quad\hat{B}=\hat{W}^{*}B,\quad\hat{C}=L_{w}^{T}P_{w}^{-1},$
which interpolates $H(s)$ at $(\mu_{1},\dots,\mu_{w})$, satisfies the following properties:
  1. 1.
    The matrix $\hat{A}$ equals $P_{w}(-S_{w}^{*}+P_{w}^{-1}\hat{B}\hat{B}^{*})P_{w}^{-1}$.
  2. 2.
    The solution $\hat{Q}_{\text{LQG}}$ to the following projected Riccati equation
    $\hat{A}^{*}\hat{Q}_{\text{LQG}}+\hat{Q}_{\text{LQG}}\hat{A}+\hat{C}^{*}\hat{C}-\hat{Q}_{\text{LQG}}\hat{B}\hat{B}^{*}\hat{Q}_{\text{LQG}}=0$ (32)
    is $P_{w}^{-1}$.
  3. 3.
    The matrix $\hat{A}-\hat{B}\hat{B}^{*}\hat{Q}_{\text{LQG}}$ is Hurwitz, with eigenvalues at $(-\mu_{1}^{*},\dots,-\mu_{w}^{*})$.
Proof.
The proof is similar to that of Theorem 3.1 and hence omitted for brevity. ∎
Again, it is clear from Theorem 3.2 that the RADI-based approximation of $Q_{\text{LQG}}$ is given by $Q_{\text{LQG}}\approx\hat{W}P_{w}^{-1}\hat{W}^{*}$.
Similar to DD-ADI-BT [2], the RADI-based non-intrusive implementation of LQG-BT can be directly obtained using the results of Theorems 3.1 and 3.2. The pseudo-code for this implementation, named “Data-driven LQG-BT (DD-LQG-BT)”, is provided in Algorithm 1.
Algorithm 1 $D$-LQG-BT
Input: RADI shifts for approximating $P_{\text{LQG}}$: $(-\sigma_{1},\cdots,-\sigma_{v})$; RADI shifts for approximating $Q_{\text{LQG}}$: $(-\mu_{1},\cdots,-\mu_{w})$; Data: $\big(H(\sigma_{1}),\cdots,H(\sigma_{v}),H(\mu_{1}),\cdots,H(\mu_{w})\big)$ and $H^{\prime}(\sigma_{i})$ for $\sigma_{i}=\mu_{j}$; Reduced order: $r$.
[…]
2: Set $S_{v}$, $L_{v}$, $S_{w}$, $L_{w}$ as in (6).
3: Compute $Q_{v}$ and $P_{w}$ by solving the Lyapunov equations (30) and (31).
4: Decompose $Q_{v}^{-1}$ and $P_{w}^{-1}$ as $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$.
5: Compute $\hat{V}_{r}$ and $\hat{W}_{r}$ from (26) and (27).
6: Compute $\hat{A}$, $\hat{B}$, and $\hat{C}$ from (28).
Remark 1.
$\mathcal{H}_{\infty}$-BT is similar to LQG-BT with the only difference being in the computation of $Q_{v}$ and $P_{w}$. In $\mathcal{H}_{\infty}$-BT, $Q_{v}$ and $P_{w}$ are computed by solving the following Lyapunov equations:
$-S_{v}^{*}Q_{v}-Q_{v}S_{v}+L_{v}^{T}L_{v}+(1-\gamma^{-2})\hat{C}^{*}\hat{C}$ $=0,$ (33)
$-S_{w}P_{w}-P_{w}S_{w}^{*}+L_{w}L_{w}^{T}+(1-\gamma^{-2})\hat{B}\hat{B}^{*}$ $=0.$ (34)
All other steps in Algorithm 1 remain the same.
3.2 RADI-based Non-intrusive Implementation of PR-BT
Let us define
$G_{\text{PR}}(s)=C_{\text{PR}}(sI-A_{\text{PR}})^{-1}B_{\text{PR}},$
where
$A_{\text{PR}}$ $=A-B_{\text{PR}}C_{\text{PR}},$ $B_{\text{PR}}$ $=BR_{\text{PR}}^{-\frac{1}{2}},$
$C_{\text{PR}}$ $=R_{\text{PR}}^{-\frac{1}{2}}C,$ $R_{\text{PR}}$ $=D+D^{T}.$
Further, define the ROM
$\hat{G}_{\text{PR}}(s)=\hat{C}_{\text{PR}}(sI-\hat{A}_{\text{PR}})^{-1}\hat{B}_{\text{PR}}$
obtained as follows:
$\hat{A}_{\text{PR}}$ $=W_{\text{PR}}^{*}(A_{\text{PR}})V_{\text{PR}}=W_{\text{PR}}^{*}(A-BR_{\text{PR}}^{-1}C)V_{\text{PR}}=\hat{A}-\hat{B}R_{\text{PR}}^{-1}\hat{C}$,
$\hat{B}_{\text{PR}}$ $=W_{\text{PR}}^{*}B_{\text{PR}}=W_{\text{PR}}^{*}BR_{\text{PR}}^{-\frac{1}{2}}=\hat{B}R_{\text{PR}}^{-\frac{1}{2}},$
$\hat{C}_{\text{PR}}$ $=C_{\text{PR}}V_{\text{PR}}=R_{\text{PR}}^{-\frac{1}{2}}CV_{\text{PR}}=R_{\text{PR}}^{-\frac{1}{2}}\hat{C}$,
where $W_{\text{PR}}^{*}V_{\text{PR}}=I$.
The Riccati equations (20) and (21) can be rewritten as follows:
$A_{\text{PR}}P_{\text{PR}}+P_{\text{PR}}A_{\text{PR}}^{T}+B_{\text{PR}}B_{\text{PR}}^{T}+P_{\text{PR}}C_{\text{PR}}^{T}C_{\text{PR}}P_{\text{PR}}=0,$ (35)
$A_{\text{PR}}^{T}Q_{\text{PR}}+Q_{\text{PR}}A_{\text{PR}}+C_{\text{PR}}^{T}C_{\text{PR}}+Q_{\text{PR}}B_{\text{PR}}B_{\text{PR}}^{T}Q_{\text{PR}}=0.$ (36)
From Theorem 3.1, the RADI-based low-rank approximation of $P_{\text{PR}}$ can be obtained as follows. The projection matrix
$V_{\text{PR}}=\begin{bmatrix}(\sigma_{1}I-A_{\text{PR}})^{-1}B_{\text{PR}}&\cdots&(\sigma_{v}I-A_{\text{PR}})^{-1}B_{\text{PR}}\end{bmatrix}$ (37)
solves the following Sylvester equation:
$A_{\text{PR}}V_{\text{PR}}-V_{\text{PR}}S_{v}+B_{\text{PR}}L_{v}=0.$ (38)
Then $\hat{C}_{\text{PR}}=C_{\text{PR}}V_{\text{PR}}$ can be computed as follows:
$\hat{C}_{\text{PR}}=\begin{bmatrix}G_{\text{PR}}(\sigma_{1})&\cdots&G_{\text{PR}}(\sigma_{v})\end{bmatrix}$. (39)
Thereafter, $Q_{v}$ is computed by solving the Lyapunov equation:
$-S_{v}^{*}Q_{v}-Q_{v}S_{v}+L_{v}^{T}L_{v}-\hat{C}_{\text{PR}}^{*}\hat{C}_{\text{PR}}=0.$ (40)
The RADI-based approximation of $P_{\text{PR}}$ is given by $V_{\text{PR}}Q_{v}^{-1}V_{\text{PR}}^{*}$.
According to Theorem 3.1, the state-space realization of the ROM $\hat{G}_{\text{PR}}(s)=\hat{C}_{\text{PR}}(sI-\hat{A}_{\text{PR}})^{-1}\hat{B}_{\text{PR}}$ (where $\hat{A}_{\text{PR}}=S_{v}-Q_{v}^{-1}L_{v}^{T}L_{v}$, $\hat{B}_{\text{PR}}=Q_{v}^{-1}L_{v}^{T}$, and $\hat{C}_{\text{PR}}=C_{\text{PR}}V_{\text{PR}}$) solves the following projected Riccati equations with $\hat{P}_{\text{PR}}=Q_{v}^{-1}>0$:
$\hat{A}_{\text{PR}}\hat{P}_{\text{PR}}+\hat{P}_{\text{PR}}\hat{A}_{\text{PR}}^{*}+\hat{B}_{\text{PR}}\hat{B}_{\text{PR}}^{*}+\hat{P}_{\text{PR}}\hat{C}_{\text{PR}}^{*}\hat{C}_{\text{PR}}\hat{P}_{\text{PR}}=0.$ (41)
Since
$\hat{A}$ $=S_{v}-Q_{v}^{-1}L_{v}^{T}L_{v}+Q_{v}^{-1}L_{v}^{T}\hat{C}_{\text{PR}},$
$\hat{B}$ $=Q_{v}^{-1}L_{v}^{T}R_{\text{PR}}^{\frac{1}{2}},\quad\hat{C}=R_{\text{PR}}^{\frac{1}{2}}\hat{C}_{\text{PR}},$ (42)
the Riccati equation (41) can be rewritten as follows:
$\hat{A}\hat{P}_{\text{PR}}+\hat{P}_{\text{PR}}\hat{A}^{*}+(\hat{B}-\hat{P}_{\text{PR}}\hat{C}^{*})R_{\text{PR}}^{-1}(\hat{B}-\hat{P}_{\text{PR}}\hat{C}^{*})^{*}=0.$ (43)
Theorem 3.3.
Let $G(s)$ be a positive-real transfer function. Define $\tilde{L}_{v}=R_{\text{PR}}^{-\frac{1}{2}}(L_{v}-\hat{C}_{\text{PR}})=\begin{bmatrix}l_{v,1}&\cdots&l_{v,v}\end{bmatrix}$ and $T_{v}=\text{blkdiag}(l_{v,1},\cdots,l_{v,v})$. Assume that:
  1. 1.
    The interpolation points $(\sigma_{1},\dots,\sigma_{v})$ are located in the right half of the $s$-plane.
  2. 2.
    The pair $(S_{v},\tilde{L}_{v})$ is observable.
  3. 3.
    The matrix $T_{v}$ is invertible.
  4. 4.
[…]
Then the following statements hold:
  1. 1.
    The projection matrix $V_{\text{PR}}$ satisfies $V_{\text{PR}}=\hat{V}T_{v}$, where $\hat{V}$ is as in (1).
  2. 2.
    The ROM $\hat{G}(s)=\hat{C}(sI-\hat{A})^{-1}\hat{B}+D$ interpolates $G(s)$ at $(\sigma_{1},\cdots,\sigma_{v})$.
  3. 3.
    The ROM $\hat{G}(s)$ is positive-real.
Proof.
By rearranging variables, the Sylvester equations (38) can be rewritten as follows:
$AV_{\text{PR}}-V_{\text{PR}}S_{v}+B\tilde{L}_{v}=0.$ (44)
It can then readily be noted that $V_{\text{PR}}=\hat{V}T_{v}$, where $\hat{V}$ is as in (1). Thus, $V_{\text{PR}}$ and $\hat{V}$ enforce the same interpolation conditions, i.e., $G(\sigma_{i})=\hat{G}(\sigma_{i})$, since the columns of $V_{\text{PR}}$ and $\hat{V}$ span the same subspace and produce the same ROM with different state-space realizations [30].
By defining $K_{\text{PR}}=(\hat{B}-\hat{P}_{\text{PR}}\hat{C}^{*})R_{\text{PR}}^{-\frac{1}{2}}$, the following holds:
$\hat{A}\hat{P}_{\text{PR}}+\hat{P}_{\text{PR}}\hat{A}^{*}$ $=-K_{\text{PR}}K_{\text{PR}}^{*}$
$\hat{P}_{\text{PR}}\hat{C}^{*}-\hat{B}$ $=-K_{\text{PR}}R_{\text{PR}}^{\frac{1}{2}}$
$R_{\text{PR}}^{\frac{1}{2}}(R_{\text{PR}}^{\frac{1}{2}})^{T}$ $=D+D^{T}.$ (45)
Since $\hat{P}_{\text{PR}}>0$, $\hat{G}(s)$ satisfies the positive-real lemma, and thus $\hat{G}(s)$ is a positive-real transfer function [34]. ∎
Now, assume that the matrices of the ROM $\hat{G}_{\text{PR}}(s)=\hat{C}_{\text{PR}}(sI-\hat{A}_{\text{PR}})^{-1}\hat{B}_{\text{PR}}$ are obtained as follows:
$\hat{A}_{\text{PR}}$ $=\hat{W}^{*}(A_{\text{PR}})\hat{V}=\hat{W}^{*}(A-BR_{\text{PR}}^{-1}C)\hat{V}=S_{v}-\zeta L_{v}-\zeta R_{\text{PR}}^{-1}C\hat{V}$,
$\hat{B}_{\text{PR}}$ $=\hat{W}^{*}B_{\text{PR}}=\hat{W}^{*}BR_{\text{PR}}^{-\frac{1}{2}}=\zeta R_{\text{PR}}^{-\frac{1}{2}},$
$\hat{C}_{\text{PR}}$ $=C_{\text{PR}}\hat{V}=R_{\text{PR}}^{-\frac{1}{2}}C\hat{V}$, (46)
where $\hat{V}$ is as in (1) and $\hat{W}$ is an unknown projection matrix satisfying $\hat{W}^{*}\hat{V}=I$. According to Theorem 3.3, the ROM $\hat{G}_{\text{PR}}(s)$ in (46) interpolates $G_{\text{PR}}(s)$ at the interpolation points $(\sigma_{1},\cdots,\sigma_{v})$ since $V_{\text{PR}}=\hat{V}T_{v}$. An immediate consequence is that a ROM $\hat{G}_{\text{PR}}(s)$ interpolating $G_{\text{PR}}(s)$ be constructed directly from the ROM $\hat{G}(s)=C\hat{V}(sI-S_{v}+\zeta L_{v})^{-1}\zeta+D$ using the formula (46). We now use this observation to derive an expression for $T_{v}$.
Note that
$\hat{C}_{\text{PR}}=C_{\text{PR}}V_{\text{PR}}=\begin{bmatrix}G_{\text{PR}}(\sigma_{1})&\cdots&G_{\text{PR}}(\sigma_{v})\end{bmatrix}=\begin{bmatrix}\hat{G}_{\text{PR}}(\sigma_{1})&\cdots&\hat{G}_{\text{PR}}(\sigma_{v})\end{bmatrix}$
$=C_{\text{PR}}\hat{V}\begin{bmatrix}(\sigma_{1}I-S_{v}+\zeta L_{v}+\zeta R_{\text{PR}}^{-1}C\hat{V})^{-1}\zeta R_{\text{PR}}^{-\frac{1}{2}}&\cdots&(\sigma_{v}I-S_{v}+\zeta L_{v}+\zeta R_{\text{PR}}^{-1}C\hat{V})^{-1}\zeta R_{\text{PR}}^{-\frac{1}{2}}\end{bmatrix}$
$=C_{\text{PR}}\hat{V}T_{v}.$
It follows directly that the matrix $T_{v}$ can be computed from the interpolant $\hat{G}(s)=C\hat{V}(sI-S_{v}+\zeta L_{v})^{-1}\zeta+D$ of $G(s)$. Such an interpolant can be generated by I-PORK, with guaranteed stability. Subsequently,
$T_{v}=\begin{bmatrix}(\sigma_{1}I-S_{v}+\zeta L_{v}+\zeta R_{\text{PR}}^{-1}C\hat{V})^{-1}\zeta R_{\text{PR}}^{-\frac{1}{2}}&\cdots&(\sigma_{v}I-S_{v}+\zeta L_{v}+\zeta R_{\text{PR}}^{-1}C\hat{V})^{-1}\zeta R_{\text{PR}}^{-\frac{1}{2}}\end{bmatrix}$
can be obtained by solving the Sylvester equation:
$(S_{v}-\zeta L_{v}-\zeta R_{\text{PR}}^{-1}C\hat{V})T_{v}-T_{v}S_{v}+\zeta R_{\text{PR}}^{-\frac{1}{2}}L_{v}=0.$ (47)
An immediate consequence is that positive-realness can be enforced non-intrusively on the ROM produced by I-PORK, using only samples of $G(s)$ at $\sigma_{i}$ and $G(\infty)$.
Next, the RADI-based approximation of $Q_{\text{PR}}$ can be obtained using Theorem 3.2 as follows. The projection matrix
$W_{\text{PR}}^{*}=\begin{bmatrix}C_{\text{PR}}(\mu_{1}I-A_{\text{PR}})^{-1}\\ \vdots\\ C_{\text{PR}}(\mu_{w}I-A_{\text{PR}})^{-1}\end{bmatrix}$ (48)
solves the Sylvester equation:
$A_{\text{PR}}^{T}W_{\text{PR}}-W_{\text{PR}}S_{w}^{*}+C_{\text{PR}}^{T}L_{w}^{T}=0.$ (49)
Then, $\hat{B}_{\text{PR}}=W_{\text{PR}}^{*}B_{\text{PR}}$ can be computed as:
$\hat{B}_{\text{PR}}=\begin{bmatrix}G_{\text{PR}}(\mu_{1})\\ \vdots\\ G_{\text{PR}}(\mu_{w})\end{bmatrix}$. (50)
Subsequently, $P_{w}$ is obtained by solving the Lyapunov equation:
$-S_{w}P_{w}-P_{w}S_{w}^{*}+L_{w}L_{w}^{T}-\hat{B}_{\text{PR}}\hat{B}_{\text{PR}}^{*}=0.$ (51)
The RADI-based approximation of $Q_{\text{PR}}$ is then given by $W_{\text{PR}}P_{w}^{-1}W_{\text{PR}}^{*}$.
According to Theorem 3.2, the state-space realization of the ROM $\hat{G}_{\text{PR}}(s)=\hat{C}_{\text{PR}}(sI-\hat{A}_{\text{PR}})^{-1}\hat{B}_{\text{PR}}$ (where $\hat{A}_{\text{PR}}=S_{w}-L_{w}L_{w}^{T}P_{w}^{-1}$, $\hat{B}_{\text{PR}}=W_{\text{PR}}^{*}B_{\text{PR}}$, and $\hat{C}_{\text{PR}}=L_{w}^{T}P_{w}^{-1}$) satisfies the projected Riccati equations with $\hat{Q}_{\text{PR}}=P_{w}^{-1}>0$
$\hat{A}_{\text{PR}}^{*}\hat{Q}_{\text{PR}}+\hat{Q}_{\text{PR}}\hat{A}_{\text{PR}}+\hat{C}_{\text{PR}}^{*}\hat{C}_{\text{PR}}+\hat{Q}_{\text{PR}}\hat{B}_{\text{PR}}\hat{B}_{\text{PR}}^{*}\hat{Q}_{\text{PR}}=0.$ (52)
Since
$\hat{A}$ $=S_{w}-L_{w}L_{w}^{T}P_{w}^{-1}+\hat{B}_{\text{PR}}L_{w}^{T}P_{w}^{-1},$
$\hat{B}$ $=\hat{B}_{\text{PR}}R_{\text{PR}}^{\frac{1}{2}},\quad\hat{C}=R_{\text{PR}}^{\frac{1}{2}}L_{w}^{T}P_{w}^{-1},$ (53)
the Riccati equation (52) can be rewritten as
$\hat{A}^{*}\hat{Q}_{\text{PR}}+\hat{Q}_{\text{PR}}\hat{A}+(\hat{C}-\hat{B}^{*}\hat{Q}_{\text{PR}})^{*}R_{\text{PR}}^{-1}(\hat{C}-\hat{B}^{*}\hat{Q}_{\text{PR}}))=0.$ (54)
Theorem 3.4.
Let $G(s)$ be a positive-real transfer function. Define $\tilde{L}_{w}^{*}=R_{\text{PR}}^{-\frac{1}{2}}(L_{w}^{T}-\hat{B}_{\text{PR}}^{*})=\begin{bmatrix}l_{w,1}^{*}&\cdots&l_{w,q}^{*}\end{bmatrix}$ and $T_{w}=\text{blkdiag}(l_{w,1}^{*},\cdots,l_{w,q}^{*})$. Assume that:
  1. 1.
    The interpolation points $(\mu_{1},\cdots,\mu_{w})$ are located in the right half of the $s$-plane.
  2. 2.
    The pair $(S_{w},\tilde{L}_{w})$ is controllable.
  3. 3.
    The matrix $T_{w}$ is invertible.
  4. 4.
[…]
Then the following statements hold:
  1. 1.
    The projection matrix $W_{\text{PR}}$ satisfies $W_{\text{PR}}=\hat{W}T_{w}$, where $\hat{W}$ is as in (2).
  2. 2.
    The ROM $\hat{G}(s)=\hat{C}(sI-\hat{A})^{-1}\hat{B}+D$ interpolates $G(s)$ at $(\mu_{1},\cdots,\mu_{w})$.
  3. 3.
    The ROM $\hat{G}(s)$ is positive-real.
Proof.
The proof is similar to the proof of Theorem 3.3 and hence omitted for brevity. ∎
We now derive an expression for $T_{w}$ when $\hat{W}$ is used to enforce interpolation conditions instead of $W_{\text{PR}}$. Since $W_{\text{PR}}$ and $\hat{W}$ enforce identical interpolation conditions, the following holds:
$\hat{B}_{\text{PR}}$ $=W_{\text{PR}}^{*}B_{\text{PR}}=\begin{bmatrix}G_{\text{PR}}(\mu_{1})\\ \vdots\\ G_{\text{PR}}(\mu_{w})\end{bmatrix}=\begin{bmatrix}\hat{G}_{\text{PR}}(\mu_{1})\\ \vdots\\ \hat{G}_{\text{PR}}(\mu_{w})\end{bmatrix}$
$=\begin{bmatrix}R_{\text{PR}}^{-\frac{1}{2}}\zeta(\mu_{1}I-S_{w}+L_{w}\zeta+\hat{W}^{*}BR_{\text{PR}}^{-1}\zeta)^{-1}\\ \vdots\\ R_{\text{PR}}^{-\frac{1}{2}}\zeta(\mu_{w}I-S_{w}+L_{w}\zeta+\hat{W}^{*}BR_{\text{PR}}^{-1}\zeta)^{-1}\end{bmatrix}\hat{W}^{*}B_{\text{PR}},$
$=T_{w}^{*}\hat{W}^{*}BR_{\text{PR}}^{-\frac{1}{2}},$
where $\hat{G}(s)=\zeta(sI-S_{w}+L_{w}\zeta)^{-1}\hat{W}^{*}B+D$ is an interpolant of $G(s)$. Such an interpolant can be generated by O-PORK, which guarantees stability, and $T_{w}$ can be constructed from this interpolant by solving the Sylvester equation:
$(S_{w}-L_{w}\zeta-\hat{W}^{*}BR_{\text{PR}}^{-1}\zeta)^{*}T_{w}-T_{w}S_{w}^{*}+\zeta^{*}R_{\text{PR}}^{-\frac{1}{2}}L_{w}^{T}=0.$
An immediate consequence is that positive-realness can be enforced non-intrusively on the ROM produced by O-PORK, using only samples of $G(s)$ at $\mu_{i}$ and $G(\infty)$.
Similar to DD-ADI-BT [2], the RADI-based non-intrusive implementation of PR-BT can be directly obtained using the results of Theorems 3.3 and 3.4. The pseudo-code for this implementation, named “Data-driven PR-BT (DD-PR-BT)”, is provided in Algorithm 2.
RADI shifts for approximating $P_{\text{PR}}$: $(-\sigma_{1},\cdots,-\sigma_{v})$; RADI shifts for approximating $Q_{\text{PR}}$: $(-\mu_{1},\cdots,-\mu_{w})$; Data: $\big(H(\sigma_{1}),\cdots,H(\sigma_{v}),H(\mu_{1}),\cdots,H(\mu_{w}),G(\infty)\big)$ and $H^{\prime}(\sigma_{i})$ for $\sigma_{i}=\mu_{j}$; Reduced order: $r$.
Output: ROM: $(\hat{A},\hat{B},\hat{C},D)$
Algorithm 2 $D$-PR-BT
[…]
2: Set $D=G(\infty)$ and $R_{\text{PR}}=D+D^{T}$.
3: Set $S_{v}$, $L_{v}$, $S_{w}$, $L_{w}$ as in (6).
4: Compute $Q_{v}$ and $P_{w}$ by solving the Lyapunov equations (9) and (10).
[…]
6: Set $\hat{C}_{\text{PR}}=R_{\text{PR}}^{-\frac{1}{2}}C\hat{V}T_{v}$ and $\hat{B}_{\text{PR}}=T_{w}^{*}\hat{W}^{*}BR_{\text{PR}}^{-\frac{1}{2}}$.
7: Update $Q_{v}$ and $P_{w}$ by solving the Lyapunov equations (40) and (51).
8: Decompose $Q_{v}^{-1}$ and $P_{w}^{-1}$ as $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$.
[…]
10: Set $W_{r}=T_{w}\hat{L}_{q}\tilde{U}_{1}\tilde{\Sigma}_{r}^{-1/2}$ and $V_{r}=T_{v}\hat{L}_{p}\tilde{V}_{1}\tilde{\Sigma}_{r}^{-1/2}$.
11: Compute $\hat{A}$, $\hat{B}$, and $\hat{C}$ from (28).
3.3 RADI-based Non-intrusive Implementation of BR-BT
By noting that $D^{T}(I-DD^{T})^{-1}=(I-D^{T}D)^{-1}D^{T}$, define the following matrices: $R_{p}=I-DD^{T}$, $R_{q}=I-D^{T}D$, $R_{b}=I+D^{T}R_{p}^{-1}D$, and $R_{c}=I+DR_{q}^{-1}D^{T}$. Further, define the transfer functions
$G_{\text{BR}}^{\text{P}}(s)$ $=C_{\text{BR}}^{\text{P}}(sI-A_{\text{BR}})^{-1}B_{\text{BR}}^{\text{P}},$
$G_{\text{BR}}^{\text{Q}}(s)$ $=C_{\text{BR}}^{\text{Q}}(sI-A_{\text{BR}})^{-1}B_{\text{BR}}^{\text{Q}},$
where
$A_{\text{BR}}$ $=A+BD^{T}R_{p}^{-1}C=A+BR_{q}^{-1}D^{T}C,$
$B_{\text{BR}}^{\text{P}}$ $=BR_{b}^{\frac{1}{2}},\quad C_{\text{BR}}^{\text{P}}=R_{p}^{-\frac{1}{2}}C,$
$B_{\text{BR}}^{\text{Q}}$ $=BR_{q}^{-\frac{1}{2}},\quad C_{\text{BR}}^{\text{Q}}=R_{c}^{\frac{1}{2}}C.$
Next, define the ROMs
$\hat{G}_{\text{BR}}^{\text{P}}(s)$ $=\hat{C}_{\text{BR}}^{\text{P}}(sI-\hat{A}_{\text{BR}})^{-1}\hat{B}_{\text{BR}}^{\text{P}},$
$\hat{G}_{\text{BR}}^{\text{Q}}(s)$ $=\hat{C}_{\text{BR}}^{\text{Q}}(sI-\hat{A}_{\text{BR}})^{-1}\hat{B}_{\text{BR}}^{\text{Q}},$
obtained as
$\hat{A}_{\text{BR}}$ $=W_{\text{BR}}^{*}A_{\text{BR}}V_{\text{BR}}=\hat{A}+\hat{B}D^{T}R_{p}^{-1}\hat{C}=\hat{A}+\hat{B}R_{q}^{-1}D^{T}\hat{C}$,
$\hat{B}_{\text{BR}}^{\text{P}}$ $=W_{\text{BR}}^{*}B_{\text{BR}}^{\text{P}}=\hat{B}R_{b}^{\frac{1}{2}},\quad\hat{C}_{\text{BR}}^{\text{P}}=C_{\text{BR}}^{\text{P}}V_{\text{BR}}=R_{p}^{-\frac{1}{2}}\hat{C}$,
$\hat{B}_{\text{BR}}^{\text{Q}}$ $=W_{\text{BR}}^{*}B_{\text{BR}}^{\text{Q}}=\hat{B}R_{q}^{-\frac{1}{2}},\quad\hat{C}_{\text{BR}}^{\text{Q}}=C_{\text{BR}}^{\text{Q}}V_{\text{BR}}=R_{c}^{\frac{1}{2}}\hat{C}$,
where $W_{\text{BR}}^{*}V_{\text{BR}}=I$.
The Riccati equations (22) and (23) can be rewritten as follows:
$A_{\text{BR}}P_{\text{BR}}+P_{\text{BR}}A_{\text{BR}}^{T}+B_{\text{BR}}^{\text{P}}(B_{\text{BR}}^{\text{P}})^{*}+P_{\text{BR}}(C_{\text{BR}}^{\text{P}})^{*}C_{\text{BR}}^{\text{P}}P_{\text{BR}}=0,$ (55)
$A_{\text{BR}}^{T}Q_{\text{BR}}+Q_{\text{BR}}A_{\text{BR}}+(C_{\text{BR}}^{\text{Q}})^{*}C_{\text{BR}}^{\text{Q}}+Q_{\text{BR}}B_{\text{BR}}^{\text{Q}}(B_{\text{BR}}^{\text{Q}})^{*}Q_{\text{BR}}=0.$ (56)
By applying Theorem 3.1, the RADI-based approximation of $P_{\text{BR}}$ can be obtained as follows. The projection matrix
$V_{\text{BR}}=\begin{bmatrix}(\sigma_{1}I-A_{\text{BR}})^{-1}B_{\text{BR}}^{\text{P}}&\cdots&(\sigma_{v}I-A_{\text{BR}})^{-1}B_{\text{BR}}^{\text{P}}\end{bmatrix}$ (57)
solves the following Sylvester equation:
$A_{\text{BR}}V_{\text{BR}}-V_{\text{BR}}S_{v}+B_{\text{BR}}^{\text{P}}L_{v}=0.$ (58)
Then, $\hat{C}_{\text{BR}}^{\text{P}}=C_{\text{BR}}^{\text{P}}V_{\text{BR}}$ can be computed as follows:
$\hat{C}_{\text{BR}}^{\text{P}}=\begin{bmatrix}G_{\text{BR}}^{\text{P}}(\sigma_{1})&\cdots&G_{\text{BR}}^{\text{P}}(\sigma_{v})\end{bmatrix}$. (59)
Thereafter, $Q_{v}$ is computed by solving the Lyapunov equation:
$-S_{v}^{*}Q_{v}-Q_{v}S_{v}+L_{v}^{T}L_{v}-(\hat{C}_{\text{BR}}^{\text{P}})^{*}\hat{C}_{\text{BR}}^{\text{P}}=0.$ (60)
The RADI-based approximation of $P_{\text{BR}}$ is given by $V_{\text{BR}}Q_{v}^{-1}V_{\text{BR}}^{*}$.
According to Theorem 3.1, the ROM $(\hat{A}_{\text{BR}}=S_{v}-\hat{B}_{\text{BR}}^{\text{P}}L_{v},\hat{B}_{\text{BR}}^{\text{P}}=Q_{v}^{-1}L_{v}^{T},\hat{C}_{\text{BR}}^{\text{P}}=C_{\text{BR}}^{\text{P}}V_{\text{BR}})$ solves the following projected Riccati equation with $\hat{P}_{\text{BR}}=Q_{v}^{-1}>0$:
$\hat{A}_{\text{BR}}\hat{P}_{\text{BR}}+\hat{P}_{\text{BR}}\hat{A}_{\text{BR}}^{*}+\hat{B}_{\text{BR}}^{\text{P}}(\hat{B}_{\text{BR}}^{\text{P}})^{*}+\hat{P}_{\text{BR}}(\hat{C}_{\text{BR}}^{\text{P}})^{*}\hat{C}_{\text{BR}}^{\text{P}}\hat{P}_{\text{BR}}=0.$ (61)
Since
$\hat{A}$ $=S_{v}-Q_{v}^{-1}L_{v}^{T}L_{v}-\hat{B}D^{T}R_{p}^{-1}\hat{C}$,
$\hat{B}$ $=Q_{v}^{-1}L_{v}^{T}R_{b}^{-\frac{1}{2}},\quad\hat{C}=R_{p}^{\frac{1}{2}}C_{\text{BR}}^{\text{P}}V_{\text{BR}},$ (62)
the Riccati equation (61) can be rewritten as
$\hat{A}P_{\text{BR}}+\hat{P}_{\text{BR}}\hat{A}^{*}+\hat{B}\hat{B}^{*}+(\hat{P}_{\text{BR}}\hat{C}^{*}+\hat{B}D^{T})R_{p}^{-1}(\hat{P}_{\text{BR}}\hat{C}^{*}+\hat{B}D^{T})^{*}=0.$
Theorem 3.5.
Let $G(s)$ be a bounded-real transfer function. Define $\tilde{L}_{v}=D^{T}R_{p}^{-\frac{1}{2}}\hat{C}_{\text{BR}}^{\text{P}}+R_{b}^{\frac{1}{2}}L_{v}=\begin{bmatrix}l_{v,1}&\cdots&l_{v,p}\end{bmatrix}$ and $T_{v}=\text{blkdiag}(l_{v,1},\cdots,l_{v,p})$. Assume that:
  1. 1.
    The interpolation points $(\sigma_{1},\dots,\sigma_{v})$ are located in the right half of the $s$-plane.
  2. 2.
    The pair $(S_{v},\tilde{L}_{v})$ is observable.
  3. 3.
    The matrix $T_{v}$ is invertible.
  4. 4.
[…]
Then the following statements hold:
  1. 1.
    The projection matrix $V_{\text{BR}}$ is equal to $\hat{V}T_{v}$ where $\hat{V}$ is as in (1).
  2. 2.
    The ROM $\hat{G}(s)=\hat{C}(sI-\hat{A})^{-1}\hat{B}+D$ interpolates $G(s)$ at $(\sigma_{1},\cdots,\sigma_{p})$.
  3. 3.
    The ROM $\hat{G}(s)$ is bounded-real.
Proof.
The proofs of 1 and 2 are similar to Theorem 3.3; hence, they are omitted for brevity.
By defining $K_{\text{BR}}=(\hat{P}_{\text{BR}}\hat{C}^{*}+\hat{B}D^{T})R_{p}^{-\frac{1}{2}}$, the following relations hold:
$\hat{A}\hat{P}_{\text{BR}}+\hat{P}_{\text{BR}}\hat{A}^{*}=-\hat{B}\hat{B}^{*}-K_{\text{BR}}K_{\text{BR}}^{*},$
$\hat{P}_{\text{BR}}\hat{C}^{*}+\hat{B}D^{T}=-K_{\text{BR}}R_{p}^{\frac{1}{2}},$
$R_{p}^{\frac{1}{2}}(R_{p}^{\frac{1}{2}})^{*}=I-DD^{T}.$
Since $\hat{P}_{\text{BR}}>0$, $\hat{G}(s)$ satisfies the bounded-real lemma [34]. ∎
We now derive an expression for $T_{v}$ when $\hat{V}$ is used to enforce interpolation conditions instead of $V_{\text{BR}}$. Since $V_{\text{BR}}$ and $\hat{V}$ enforce the same interpolation conditions, the following holds:
$\hat{C}_{\text{BR}}^{\text{P}}$ $=C_{\text{BR}}^{\text{P}}V_{\text{BR}}=\begin{bmatrix}G_{\text{BR}}^{\text{P}}(\sigma_{1})&\cdots&G_{\text{BR}}^{\text{P}}(\sigma_{v})\end{bmatrix}=\begin{bmatrix}\hat{G}_{\text{BR}}^{\text{P}}(\sigma_{1})&\cdots&\hat{G}_{\text{BR}}^{\text{P}}(\sigma_{v})\end{bmatrix}$
$=C_{\text{BR}}^{\text{P}}\hat{V}\begin{bmatrix}(\sigma_{1}I-S_{v}+\zeta L_{v}-\zeta D^{T}R_{p}^{-1}C\hat{V})^{-1}\zeta R_{b}^{\frac{1}{2}}&\cdots&(\sigma_{v}I-S_{v}+\zeta L_{v}-\zeta D^{T}R_{p}^{-1}C\hat{V})^{-1}\zeta R_{b}^{\frac{1}{2}}\end{bmatrix}$
$=C_{\text{BR}}^{\text{P}}\hat{V}T_{v}=R_{p}^{-\frac{1}{2}}C\hat{V}T_{v},$
where $\hat{G}(s)=C\hat{V}(sI-S_{v}+\xi L_{v})^{-1}\xi+D$ is an interpolant of $G(s)$. Such an interpolant can be produced by I-PORK, which is guaranteed to be stable, and $T_{v}$ can be constructed from this interpolant. Specifically, $T_{v}$ is the solution to the following Sylvester equation:
$(S_{v}-\zeta L_{v}+\zeta D^{T}R_{p}^{-1}C\hat{V})T_{v}-T_{v}S_{v}+\zeta R_{b}^{\frac{1}{2}}L_{v}=0.$
An immediate consequence of this observation is that bounded-realness can be enforced non-intrusively on the ROM produced by I-PORK using only samples of $G(s)$ at $\sigma_{i}$ and the sample $G(\infty)$.
As per Theorem 3.2, the RADI-based approximation of $Q_{\text{BR}}$ can be obtained as follows. The projection matrix
$W_{\text{BR}}^{*}=\begin{bmatrix}C_{\text{BR}}^{\text{Q}}(\mu_{1}I-A_{\text{BR}})^{-1}\\ \vdots\\ C_{\text{BR}}^{\text{Q}}(\mu_{w}I-A_{\text{BR}})^{-1}\end{bmatrix}$ (63)
solves the following Sylvester equation:
$A_{\text{BR}}^{T}W_{\text{BR}}-W_{\text{BR}}S_{w}^{*}+(C_{\text{BR}}^{\text{Q}})^{*}L_{w}^{T}=0.$ (64)
Then, $\hat{B}_{\text{BR}}^{\text{Q}}=W_{\text{BR}}^{*}B_{\text{BR}}^{\text{Q}}$ can be computed as:
$\hat{B}_{\text{BR}}^{\text{Q}}=\begin{bmatrix}G_{\text{BR}}^{\text{Q}}(\mu_{1})\\ \vdots\\ G_{\text{BR}}^{\text{Q}}(\mu_{w})\end{bmatrix}$. (65)
Thereafter, $P_{w}$ is computed by solving the Lyapunov equation:
$-S_{w}P_{w}-P_{w}S_{w}^{*}+L_{w}L_{w}^{T}-\hat{B}_{\text{BR}}^{\text{Q}}(\hat{B}_{\text{BR}}^{\text{Q}})^{*}=0.$ (66)
The RADI-based approximation of $Q_{\text{BR}}$ is given by $W_{\text{BR}}P_{w}^{-1}W_{\text{BR}}^{*}$.
According to Theorem 3.2, the ROM $(\hat{A}_{\text{BR}}=S_{w}-L_{w}\hat{C}_{\text{BR}}^{\text{Q}},\hat{B}_{\text{BR}}^{\text{Q}}=W_{\text{BR}}^{*}B_{\text{BR}}^{\text{Q}},\hat{C}_{\text{BR}}^{\text{Q}}=L_{w}^{T}P_{w}^{-1})$ solves the following projected Riccati equation with $\hat{Q}_{\text{BR}}=P_{w}^{-1}>0$:
$\hat{A}_{\text{BR}}^{*}\hat{Q}_{\text{BR}}+\hat{Q}_{\text{BR}}\hat{A}_{\text{BR}}+(\hat{C}_{\text{BR}}^{\text{Q}})^{*}\hat{C}_{\text{BR}}^{\text{Q}}+\hat{Q}_{\text{BR}}\hat{B}_{\text{BR}}^{\text{Q}}(\hat{B}_{\text{BR}}^{\text{Q}})^{*}Q_{\text{BR}}=0.$ (67)
Since
$\hat{A}$ $=S_{w}-L_{w}L_{w}^{T}P_{w}^{-1}-\hat{B}R_{q}^{-1}D^{T}\hat{C}$,
$\hat{B}$ $=\hat{W}_{\text{BR}}^{*}B_{\text{BR}}^{\text{Q}}R_{q}^{\frac{1}{2}},\quad\hat{C}=R_{c}^{-\frac{1}{2}}L_{w}^{T}P_{w}^{-1},$ (68)
the Riccati equation (67) can be rewritten as:
$\hat{A}^{*}\hat{Q}_{\text{BR}}+\hat{Q}_{\text{BR}}\hat{A}+\hat{C}^{*}\hat{C}+(\hat{B}^{*}\hat{Q}_{\text{BR}}+D^{T}\hat{C})^{*}R_{q}^{-1}(\hat{B}^{*}\hat{Q}_{\text{BR}}+D^{T}\hat{C})=0.$ (69)
Theorem 3.6.
Let $G(s)$ be a bounded-real transfer function. Define $\tilde{L}_{w}^{*}=D^{T}R_{p}^{-\frac{1}{2}}\hat{B}_{\text{BR}}^{\text{Q}}+R_{c}^{\frac{1}{2}}L_{w}^{T}=\begin{bmatrix}l_{w,1}^{*}&\cdots&l_{w,q}^{*}\end{bmatrix}$ and $T_{w}=\text{blkdiag}(l_{w,1}^{*},\cdots,l_{w,q}^{*})$. Assume that:
  1. 1.
    The interpolation points $(\mu_{1},\cdots,\mu_{w})$ are located in the right half of the $s$-plane.
  2. 2.
    The pair $(S_{w},\tilde{L}_{w})$ is controllable.
  3. 3.
    The matrix $T_{w}$ is invertible.
  4. 4.
[…]
Then the following statements hold:
  1. 1.
    The projection matrix $W_{\text{BR}}$ is equal to $\hat{W}T_{w}$ where $\hat{W}$ is as in (2).
  2. 2.
    The ROM $\hat{G}(s)=\hat{C}(sI-\hat{A})^{-1}\hat{B}+D$ interpolates $G(s)$ at $(\mu_{1},\cdots,\mu_{q})$.
  3. 3.
    The ROM $\hat{G}(s)$ is bounded-real.
Proof.
The proof is similar to that of Theorem 3.5 and hence omitted for brevity. ∎
We now derive an expression for $T_{w}$ when $\hat{W}$ is used to enforce interpolation conditions instead of $W_{\text{BR}}$. Since $W_{\text{BR}}$ and $\hat{W}$ enforce identical interpolation conditions, the following holds:
$\hat{B}_{\text{BR}}^{\text{Q}}$ $=W_{\text{BR}}^{*}B_{\text{BR}}^{\text{Q}}=\begin{bmatrix}G_{\text{BR}}^{\text{Q}}(\mu_{1})\\ \vdots\\ G_{\text{BR}}^{\text{Q}}(\mu_{w})\end{bmatrix}=\begin{bmatrix}\hat{G}_{\text{PR}}^{\text{Q}}(\mu_{1})\\ \vdots\\ \hat{G}_{\text{PR}}^{\text{Q}}(\mu_{w})\end{bmatrix}$
$=\begin{bmatrix}R_{c}^{\frac{1}{2}}\zeta(\mu_{1}I-S_{w}+L_{w}\zeta-\hat{W}^{*}BR_{q}^{-1}D^{T}\zeta)^{-1}\\ \vdots\\ R_{c}^{\frac{1}{2}}\zeta(\mu_{w}I-S_{w}+L_{w}\zeta-\hat{W}^{*}BR_{q}^{-1}D^{T}\zeta)^{-1}\end{bmatrix}\hat{W}^{*}B_{\text{BR}}^{\text{Q}},$
$=T_{w}^{*}\hat{W}^{*}BR_{q}^{-\frac{1}{2}},$
where $\hat{G}(s)=\zeta(sI-S_{w}+L_{w}\zeta)^{-1}\hat{W}^{*}B+D$ is an interpolant of $G(s)$. This interpolant can be produced by O-PORK, which is guaranteed to be stable, and $T_{w}$ can be constructed from it. Specifically, $T_{w}$ is obtained by solving the Sylvester equation:
$(S_{w}-L_{w}\zeta+\hat{W}^{*}BR_{q}^{-1}D^{T}\zeta)^{*}T_{w}-T_{w}S_{w}^{*}+\zeta^{*}R_{c}^{\frac{1}{2}}L_{w}^{T}=0.$ (70)
An immediate consequence is that bounded-realness can be enforced non-intrusively on the ROM produced by O-PORK, using only samples of $G(s)$ at $\mu_{i}$ and $G(\infty)$.
Similar to DD-ADI-BT [2], the RADI-based non-intrusive implementation of BR-BT can be directly obtained using the results of Theorems 3.5 and 3.6. The pseudo-code for this implementation, named “Data-driven BR-BT (DD-BR-BT)”, is provided in Algorithm 3.
RADI shifts for approximating $P_{\text{BR}}$: $(-\sigma_{1},\cdots,-\sigma_{v})$; RADI shifts for approximating $Q_{\text{BR}}$: $(-\mu_{1},\cdots,-\mu_{w})$; Data: $\big(H(\sigma_{1}),\cdots,H(\sigma_{v}),H(\mu_{1}),\cdots,H(\mu_{w}),G(\infty)\big)$ and $H^{\prime}(\sigma_{i})$ for $\sigma_{i}=\mu_{j}$; Reduced order: $r$.
Output: ROM: $(\hat{A},\hat{B},\hat{C},D)$
Algorithm 3 $D$-BR-BT
[…]
2: Set $D=G(\infty)$, $R_{p}=I-DD^{T}$, $R_{b}=I+D^{T}R_{p}^{-1}D$, $R_{q}=I-D^{T}D$, and $R_{c}=I+DR_{q}^{-1}D^{T}$.
3: Set $S_{v}$, $L_{v}$, $S_{w}$, and $L_{w}$ as in (6).
4: Compute $Q_{v}$ and $P_{w}$ by solving the Lyapunov equations (9) and (10).
[…]
6: Set $\hat{C}_{\text{BR}}^{\text{P}}=R_{p}^{-\frac{1}{2}}C\hat{V}T_{v}$ and $\hat{B}_{\text{BR}}^{\text{Q}}=T_{w}^{*}\hat{W}^{*}BR_{q}^{-\frac{1}{2}}$.
7: Update $Q_{v}$ and $P_{w}$ by solving the Lyapunov equations (60) and (66).
8: Decompose $Q_{v}^{-1}$ and $P_{w}^{-1}$ as $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$.
[…]
10: Set $W_{r}=T_{w}\hat{L}_{q}\tilde{U}_{1}\tilde{\Sigma}_{r}^{-1/2}$ and $V_{r}=T_{v}\hat{L}_{p}\tilde{V}_{1}\tilde{\Sigma}_{r}^{-1/2}$.
11: Compute $\hat{A}$, $\hat{B}$, and $\hat{C}$ from (28).
3.4 ADI-based Non-intrusive Implementation of SW-BT
Let us define the transfer function
$G_{\text{SW}}(s)=C_{\text{SW}}(sI-A_{\text{SW}})^{-1}B,$
where $A_{\text{SW}}=A-BC_{\text{SW}}$ and $C_{\text{SW}}=D^{-1}C$. Next, define the ROM $\hat{G}_{\text{SW}}(s)$
$\hat{G}_{\text{SW}}(s)=\hat{C}_{\text{SW}}(sI-\hat{A}_{\text{SW}})^{-1}\hat{B}$,
obtained as
$\hat{A}_{\text{SW}}$ $=W_{\text{SW}}^{*}A_{\text{SW}}\hat{V}=\hat{A}-\hat{B}D^{-1}\hat{C}$,
$\hat{B}$ $=W_{\text{SW}}^{*}B,\quad\hat{C}_{\text{SW}}=C_{\text{SW}}V_{\text{SW}}=D^{-1}\hat{C}$,
where $W_{\text{SW}}^{*}V_{\text{SW}}=I$.
The Lyapunov equation (24) can be rewritten as:
$A_{\text{SW}}^{T}Q_{\text{SW}}+Q_{\text{SW}}A_{\text{SW}}+C_{\text{SW}}^{T}C_{\text{SW}}=0.$ (71)
The O-PORK-based approximation of $Q_{\text{SW}}$ (equivalent to the ADI method) is computed as follows. First, solve the Sylvester equation for the projection matrix $W_{\text{SW}}$:
$A_{\text{SW}}^{T}W_{\text{SW}}-W_{\text{SW}}S_{w}^{*}+C_{\text{SW}}^{T}L_{w}^{T}=0.$ (72)
Then, compute $P_{w}$ by solving the Lyapunov equation (10). The ADI-based approximation of $Q_{\text{SW}}$ is obtained as: $Q_{\text{S}}\approx W_{\text{SW}}P_{w}^{-1}W_{\text{SW}}^{*}$. The following ROM $\hat{G}(s)$ can be recovered from the ROM $\hat{G}_{\text{SW}}(s)=L_{w}^{T}P_{w}^{-1}(sI-S_{w}+L_{w}L_{w}^{T}P_{w}^{-1})^{-1}W_{\text{SW}}^{*}B$ produced by O-PORK:
$\hat{A}$ $=S_{w}+L_{w}L_{w}^{T}P_{w}^{-1}+W_{\text{S}}^{*}BL_{w}^{T}P_{w}^{-1},$
$\hat{B}$ $=W_{\text{SW}}^{*}B,\quad\hat{C}=DL_{w}^{T}P_{w}^{-1}.$ (73)
Theorem 3.7.
Let $G(s)$ be a square stable minimum phase transfer function. Define $\tilde{L}_{w}^{*}=D^{-T}(L_{w}^{T}-\hat{B}^{*})=\begin{bmatrix}l_{w,1}^{*}&\cdots&l_{w,q}^{*}\end{bmatrix}$ and $T_{w}=\text{blkdiag}(l_{w,1}^{*},\cdots,l_{w,q}^{*})$. Assume that:
  1. 1.
    The interpolation points $(\mu_{1},\cdots,\mu_{w})$ are located in the right half of the $s$-plane.
  2. 2.
    The pair $(S_{w},\tilde{L}_{w})$ is controllable.
  3. 3.
    The matrix $T_{w}$ is invertible.
  4. 4.
[…]
Then the following statements hold:
  1. 1.
    The projection matrix $W_{\text{SW}}$ is equal to $\hat{W}T_{w}$ where $\hat{W}$ is as in (2).
  2. 2.
    The ROM $\hat{G}(s)=\hat{C}(sI-\hat{A})^{-1}\hat{B}+D$ interpolates $G(s)$ at $(\mu_{1},\cdots,\mu_{q})$.
  3. 3.
    The ROM $\hat{G}(s)$ is minimum phase.
Proof.
The proof of 1 and 2 is similar to that of Theorem 3.4 and hence omitted for brevity.
The zeros of $\hat{G}(s)$, i.e., the eigenvalues of $\hat{A}_{\text{SW}}=\hat{A}-\hat{B}D^{-1}\hat{C}$, are $(-\mu_{1}^{*},\cdots,-\mu_{w}^{*})$ in O-PORK. Since the interpolation points in O-PORK are in the right half of the $s$-plane, $\hat{A}_{\text{SW}}$ is Hurwitz and $\hat{G}(s)$ is minimum phase. ∎
We now derive an expression for $T_{w}$ when $\hat{W}$ is used for enforcing interpolation conditions instead of $W_{\text{SW}}$. Since $W_{\text{SW}}$ and $\hat{W}$ enforce the same interpolation conditions, the following holds:
$\hat{B}$ $=W_{\text{SW}}^{*}B=\begin{bmatrix}G_{\text{SW}}(\mu_{1})\\ \vdots\\ G_{\text{SW}}(\mu_{w})\end{bmatrix}=\begin{bmatrix}\hat{G}_{\text{SW}}(\mu_{1})\\ \vdots\\ \hat{G}_{\text{SW}}(\mu_{w})\end{bmatrix}$
$=\begin{bmatrix}D^{-1}\zeta(\mu_{1}I-S_{w}+L_{w}\zeta+\hat{W}^{*}BD^{-1}\zeta)^{-1}\\ \vdots\\ D^{-1}\zeta(\mu_{w}I-S_{w}+L_{w}\zeta+\hat{W}^{*}BD^{-1}\zeta)^{-1}\end{bmatrix}\hat{W}^{*}B,$
$=T_{w}^{*}\hat{W}^{*}B,$
where $\hat{G}(s)=\zeta(sI-S_{w}+L_{w}\zeta)^{-1}\hat{W}^{*}B+D$ is the interpolant of $G(s)$. Such an interpolant can be produced by O-PORK, which is guaranteed to be stable, and $T_{w}$ can be constructed from this interpolant. $T_{w}$ can be computed by solving the following Sylvester equation:
$\big(S_{w}-L_{w}\zeta-\hat{W}^{*}BD^{-1}\zeta\big)^{*}T_{w}-T_{w}S_{w}^{*}+\zeta^{*}D^{-T}L_{w}^{T}=0.$
An immediate consequence of this observation is that the minimum-phase property can be enforced on the ROM nonintrusively produced by O-PORK using only samples of $G(s)$ at $\mu_{i}$ and the sample $G(\infty)$.
Similar to DD-ADI-BT [2], the ADI-based nonintrusive implementation of SW-BT can be derived directly using the results of Theorem 3.7. The pseudo-code for this implementation, named “Data-driven SW-BT (DD-SW-BT)”, is provided in Algorithm 4.
ADI shifts for approximating $P$: $(-\sigma_{1},\cdots,-\sigma_{v})$; ADI shifts for approximating $Q_{\text{SW}}$: $(-\mu_{1},\cdots,-\mu_{w})$; Data: $\big(H(\sigma_{1}),\cdots,H(\sigma_{v}),H(\mu_{1}),\cdots,H(\mu_{w}),G(\infty)\big)$ and $H^{\prime}(\sigma_{i})$ for $\sigma_{i}=\mu_{j}$; Reduced order: $r$.
Output: ROM: $(\hat{A},\hat{B},\hat{C},D)$
Algorithm 4 $D$-SW-BT
[…]
2: Set $S_{v}$, $L_{v}$, $S_{w}$, and $L_{w}$ as in (6).
3: Compute $Q_{v}$ and $P_{w}$ by solving the Lyapunov equations (9) and (10).
[…]
5: Decompose $Q_{v}^{-1}$ and $P_{w}^{-1}$ as $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$.
[…]
7: Set $W_{r}=T_{w}\hat{L}_{q}\tilde{U}_{1}\tilde{\Sigma}_{r}^{-1/2}$ and $V_{r}=\hat{L}_{p}\tilde{V}_{1}\tilde{\Sigma}_{r}^{-1/2}$.
8: Compute $\hat{A}$, $\hat{B}$, and $\hat{C}$ from (28).
3.5 ADI/RADI-based Non-intrusive Implementation of BST
Let us define the transfer function
$G_{\text{S}}(s)=C_{\text{S}}(sI-A_{\text{S}})^{-1}B_{\text{S}},$
where
$A_{\text{S}}$ $=A-B_{\text{S}}C_{\text{S}},$ $B_{\text{S}}$ $=(PC^{T}+BD^{T})R_{\text{S}}^{-\frac{1}{2}},$
$C_{\text{S}}$ $=R_{\text{S}}^{-\frac{1}{2}}C,$ $R_{\text{S}}$ $=DD^{T}.$
The Riccati equation (25) can be rewritten as follows:
$A_{\text{S}}^{T}Q_{\text{S}}+Q_{\text{S}}A_{\text{S}}+C_{\text{S}}^{T}C_{\text{S}}+Q_{\text{S}}B_{\text{S}}B_{\text{S}}^{T}Q_{\text{S}}=0.$ (74)
Unlike the Riccati equations considered up until now, the RADI-based approximation of this Riccati equation is a bit more involved. We first need to replace $B_{\text{S}}$ with its approximation. Let us replace $P$ in $B_{\text{S}}$ with $\hat{V}Q_{v}^{-1}\hat{V}^{*}$, which is an ADI-based approximation of $P$ obtained via I-PORK. Then, $B_{\text{S}}\approx\tilde{B}_{\text{S}}=(\hat{V}Q_{v}^{-1}\hat{V}^{*}C^{T}+BD^{T})R_{\text{S}}^{-\frac{1}{2}}$ and $A_{\text{S}}\approx\tilde{A}_{\text{S}}=A-\tilde{B}_{\text{S}}C_{\text{S}}$. Thereafter, we proceed with the RADI-based approximation of the following Riccati equation:
$\tilde{A}_{\text{S}}^{*}\tilde{Q}_{\text{S}}+\tilde{Q}_{\text{S}}\tilde{A}_{\text{S}}+C_{\text{S}}^{*}C_{\text{S}}+\tilde{Q}_{\text{S}}\tilde{B}_{\text{S}}\tilde{B}_{\text{S}}^{*}\tilde{Q}_{\text{S}}=0.$ (75)
The RADI-based approximation of $\tilde{Q}_{\text{S}}$ can be obtained using Theorem 3.2 as follows. The projection matrix $W_{\text{S}}$ is computed by solving the following Sylvester equation:
$\tilde{A}_{\text{S}}^{*}W_{\text{S}}-W_{\text{S}}S_{w}^{*}+C_{\text{S}}^{T}L_{w}^{T}=0.$ (76)
The matrix $P_{w}$ is computed by solving the following Sylvester equation:
$-S_{w}P_{w}-P_{w}S_{w}^{*}+L_{w}L_{w}^{T}-\hat{B}_{\text{S}}\hat{B}_{\text{S}}^{*}=0,$ (77)
where $\hat{B}_{\text{S}}=W_{\text{S}}^{*}\tilde{B}_{\text{S}}$. The RADI-based approximation of $\tilde{Q}_{\text{S}}$ is given by $W_{\text{S}}P_{w}^{-1}W_{\text{S}}^{*}$.
Once again, it can be shown that $W_{\text{S}}$ and $\hat{W}$ in (2) enforce the same interpolation conditions and $W_{\text{S}}=\hat{W}T_{w}$, where $T_{w}$ is invertible. The proof is similar and hence omitted here for brevity.
Let us define the ROM
$\hat{G}_{\text{S}}(s)=\hat{C}_{\text{S}}(sI-\hat{A}_{\text{S}})^{-1}\hat{B}_{\text{S}},$
obtained as
$\hat{A}_{\text{S}}=W_{\text{\text{S}}}^{*}\tilde{A}_{\text{S}}V_{\text{S}},\quad\hat{B}_{\text{S}}=W_{\text{\text{S}}}^{*}\tilde{B}_{\text{S}},\quad\hat{C}_{\text{S}}=C_{\text{S}}V_{\text{S}},$
where $W_{\text{\text{S}}}^{*}V_{\text{S}}=I$.
We now derive an expression for $T_{w}$ when $\hat{W}$ is used for enforcing interpolation conditions instead of $W_{\text{S}}$. Since $W_{\text{S}}$ and $\hat{W}$ enforce the same interpolation conditions, the following holds:
$\hat{B}_{\text{S}}$ $=W_{\text{S}}^{*}\tilde{B}_{\text{S}}=\begin{bmatrix}G_{\text{S}}(\mu_{1})\\ \vdots\\ G_{\text{S}}(\mu_{w})\end{bmatrix}=\begin{bmatrix}\hat{G}_{\text{S}}(\mu_{1})\\ \vdots\\ \hat{G}_{\text{S}}(\mu_{w})\end{bmatrix}$
$=\begin{bmatrix}R_{\text{S}}^{-\frac{1}{2}}\zeta\big(\mu_{1}I-S_{w}+L_{w}\zeta+(\hat{W}^{*}\hat{V}Q_{v}^{-1}\hat{V}^{*}C^{T}+\hat{W}^{*}BD^{T})R_{\text{S}}^{-1}\zeta\big)^{-1}\\ \vdots\\ R_{\text{S}}^{-\frac{1}{2}}\zeta\big(\mu_{w}I-S_{w}+L_{w}\zeta+(\hat{W}^{*}\hat{V}Q_{v}^{-1}\hat{V}^{*}C^{T}+\hat{W}*BD^{T})R_{\text{S}}^{-1}\zeta\big)^{-1}\end{bmatrix}\hat{W}^{*}B_{\text{X}},$
$=T_{w}^{*}\hat{W}^{*}B_{\text{S}},$
$=T_{w}^{*}\big((\hat{W}^{*}\hat{V})Q_{v}^{-1}(C\hat{V})^{*}+(\hat{W}^{*}B)D^{T}\big)R_{\text{S}}^{-\frac{1}{2}},$
where $\hat{G}(s)=\zeta(sI-S_{w}+L_{w}\zeta)^{-1}\hat{W}^{*}B+D$ is the interpolant of $G(s)$. Such an interpolant can be produced by O-PORK, which is guaranteed to be stable. $T_{w}$ can be constructed from this interpolant by solving the following Sylvester equation:
$\big(S_{w}-L_{w}\zeta-[(\hat{W}^{*}\hat{V})Q_{v}^{-1}(C\hat{V})^{*}+(\hat{W}^{*}B)D^{T}]$ $R_{\text{S}}^{-1}\zeta\big)^{*}T_{w}$
$-T_{w}S_{w}^{*}+\zeta^{*}R_{\text{S}}^{-\frac{1}{2}}L_{w}^{T}=0.$
Similar to DD-ADI-BT [2], the ADI- and RADI-based nonintrusive implementation of BST can be obtained readily. The pseudo-code for this implementation, named “Data-driven BST (DD-BST)”, is provided in Algorithm 5.
ADI shifts for approximating $P$: $(-\sigma_{1},\cdots,-\sigma_{v})$; RADI shifts for approximating $Q_{\text{S}}$: $(-\mu_{1},\cdots,-\mu_{w})$; Data: $\big(H(\sigma_{1}),\cdots,H(\sigma_{v}),H(\mu_{1}),\cdots,H(\mu_{w}),G(\infty)\big)$ and $H^{\prime}(\sigma_{i})$ for $\sigma_{i}=\mu_{j}$; Reduced order: $r$.
Output: ROM: $(\hat{A},\hat{B},\hat{C},D)$
Algorithm 5 $D$-BST
[…]
2: Set $D=G(\infty)$ and $R_{\text{S}}=DD^{T}$.
3: Set $S_{v}$, $L_{v}$, $S_{w}$, and $L_{w}$ as in (6).
4: Compute $Q_{v}$ and $P_{w}$ by solving the Lyapunov equations (9) and (10).
[…]
6: Set $\hat{B}_{\text{S}}=T_{w}^{*}\big((\hat{W}^{*}\hat{V})Q_{v}^{-1}(C\hat{V})^{*}+(\hat{W}^{*}B)D^{T}\big)R_{\text{S}}^{-\frac{1}{2}}$.
7: Update $P_{w}$ by solving the Sylvester equation (77).
8: Decompose $Q_{v}^{-1}$ and $P_{w}^{-1}$ as $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$.
[…]
10: Set $W_{r}=T_{w}\hat{L}_{q}\tilde{U}_{1}\tilde{\Sigma}_{r}^{-1/2}$ and $V_{r}=\hat{L}_{p}\tilde{V}_{1}\tilde{\Sigma}_{r}^{-1/2}$.
11: Compute $\hat{A}$, $\hat{B}$, and $\hat{C}$ from (28).
4 Practical Considerations in Implementation
The theoretical foundations of the ADI [16] and RADI [4] methods are well established. Their performance in approximating Lyapunov and Riccati equations is also well documented. However, the implementation presented in this paper involves inverses $Q_{v}^{-1}$ and $P_{w}^{-1}$, which can cause numerical problems as the number of samples used for implementation increases. This is because the solution of Lyapunov equations is generally low-rank, enabling low-rank approximation via the low-rank ADI method [16] in the first place. Thus, as the number of samples used for implementation increases, the matrices $Q_{v}$ and $P_{w}$ start losing rank. Note that these matrices do not appear in the original algorithms. In data-driven modeling, one typically uses as many samples as available to ensure maximum accuracy. As discussed in [2], the inverses $Q_{v}^{-1}$ and $P_{w}^{-1}$ in DD-ADI-BT can be computed conveniently when the interpolation points are lightly damped, i.e.,
$\sigma_{i}=\Bigg(\frac{\zeta_{i,\sigma}|\omega_{i,\sigma}|}{\sqrt{1-\zeta_{i,\sigma}^{2}}}\Bigg)+j\omega_{i,\sigma}\quad\textnormal{and}\quad\mu_{i}=\Bigg(\frac{\zeta_{i,\mu}|\omega_{i,\mu}|}{\sqrt{1-\zeta_{i,\mu}^{2}}}\Bigg)+j\omega_{i,\mu},$ (78)
wherein $\zeta_{i,\sigma}\ll 1$ and $\zeta_{i,\mu}\ll 1$. If we select the damping coefficients $\zeta_{i,\sigma}\approx 0$ and $\zeta_{i,\mu}\approx 0$, this means we are sampling $G(s)$ in close proximity to the $j\omega$-axis, which is also practically reasonable for capturing the frequency response of $G(s)$. When the interpolation points are lightly damped (i.e., $\zeta_{i,\sigma}\ll 1$ and $\zeta_{i,\mu}\ll 1$), the respective matrices $Q_{v}$, $P_{w}$, $Q_{v}^{-1}$, $P_{w}^{-1}$, $T_{v}$, and $T_{w}$ become diagonally dominant due to the modal form of the matrices $S_{v}$ and $S_{w}$ that appear in their associated Lyapunov, Sylvester, and Riccati equations. If $\zeta_{i,\sigma}\approx 0$ and $\zeta_{i,\mu}\approx 0$, the off-diagonal elements are nearly zero. For details and analytical formulas for computing the diagonal elements of the Lyapunov, Sylvester, and Riccati equations in this case, see [40]. Essentially, this means that each diagonal element of $Q_{v}^{-1}$ and $P_{w}^{-1}$ depends solely on its respective pair $(\sigma_{i},G(\sigma_{i}))$ or $(\mu_{i},G(\mu_{i}))$, while the dependence on other samples or sampling points becomes negligible.
Assuming that the interpolation points are lightly damped, the expressions for the diagonal entries of $Q_{v}^{-1}$ and $P_{w}^{-1}$ can be derived using the analytical formulas in [40] as follows:
For DD-LQG-BT:
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\sigma_{1})[I_{m}+H^{*}(\sigma_{1})H(\sigma_{1})]^{-1},\cdots,$
$\hskip 113.81102pt2\text{Real}(\sigma_{v})[I_{m}+H^{*}(\sigma_{v})H(\sigma_{v})]^{-1}\big),$ (79)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\mu_{1})[I_{p}+H(\mu_{1})H^{*}(\mu_{1})]^{-1},\cdots,$
$\hskip 113.81102pt2\text{Real}(\mu_{w})[I_{p}+H(\mu_{w})H^{*}(\mu_{w})]^{-1}\big).$ (80)
For $D$-$\mathcal{H}_{\infty}$-BT:
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\sigma_{1})[I_{m}+(1-\gamma^{-2})H^{*}(\sigma_{1})H(\sigma_{1})]^{-1},\cdots,$
$\hskip 85.35826pt2\text{Real}(\sigma_{v})[I_{m}+(1-\gamma^{-2})H^{*}(\sigma_{v})H(\sigma_{v})]^{-1}\big),$ (81)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\mu_{1})[I_{p}+(1-\gamma^{-2})H(\mu_{1})H^{*}(\mu_{1})]^{-1},\cdots,$
$\hskip 85.35826pt2\text{Real}(\mu_{w})[I_{p}+(1-\gamma^{-2})H(\mu_{w})H^{*}(\mu_{w})]^{-1}\big).$ (82)
For DD-PR-BT:
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\sigma_{1})[I_{m}-t_{v,1}^{*}H^{*}(\sigma_{1})R_{\text{PR}}^{-1}H(\sigma_{1})t_{v,1}]^{-1},\cdots,$
$\hskip 56.9055pt2\text{Real}(\sigma_{v})[I_{m}-t_{v,v}^{*}H^{*}(\sigma_{v})R_{\text{PR}}^{-1}H(\sigma_{v})t_{v,v}]^{-1}\big),$ (83)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\mu_{1})[I_{p}-t_{w,1}^{*}H(\mu_{1})R_{\text{PR}}^{-1}H^{*}(\mu_{1})t_{w,1}]^{-1},\cdots,$
$\hskip 56.9055pt2\text{Real}(\mu_{w})[I_{p}-t_{w,w}^{*}H(\mu_{w})R_{\text{PR}}^{-1}H^{*}(\mu_{w})t_{w,w}]^{-1}\big),$ (84)
$T_{v}$ $\approx\text{blkdiag}\big(t_{v,1},\cdots,t_{v,v}\big),$ (85)
$T_{w}$ $\approx\text{blkdiag}\big(t_{w,1},\cdots,t_{v,w}\big),$ (86)
where
$t_{v,i}$ $=[I_{m}+R_{\text{PR}}^{-1}H(\sigma_{i})]^{-1}R_{\text{PR}}^{-\frac{1}{2}},$ (87)
$t_{w,i}$ $=[I_{m}+R_{\text{PR}}^{-1}H^{*}(\mu_{i})]^{-1}R_{\text{PR}}^{-\frac{1}{2}}.$ (88)
For DD-BR-BT:
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\sigma_{1})[I_{m}-t_{v,1}^{*}H^{*}(\sigma_{1})R_{p}^{-1}H(\sigma_{1})t_{v,1}]^{-1},\cdots,$
$\hskip 56.9055pt2\text{Real}(\sigma_{v})[I_{m}+t_{v,v}^{*}H^{*}(\sigma_{v})R_{p}^{-1}H(\sigma_{v})t_{v,v}]^{-1}\big),$ (89)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\mu_{1})[I_{p}-t_{w,1}^{*}H(\mu_{1})R_{q}^{-1}H^{*}(\mu_{1})t_{w,1}]^{-1},\cdots,$
$\hskip 56.9055pt2\text{Real}(\mu_{w})[I_{p}-t_{w,w}^{*}H(\mu_{w})R_{q}^{-1}H^{*}(\mu_{w})t_{w,w}]^{-1}\big),$ (90)
$T_{v}$ $\approx\text{blkdiag}\big(t_{v,1},\cdots,t_{v,v}\big),$ (91)
$T_{w}$ $\approx\text{blkdiag}\big(t_{w,1},\cdots,t_{v,w}\big),$ (92)
where
$t_{v,i}$ $=[I_{m}-D^{T}R_{p}^{-1}H(\sigma_{i})]^{-1}R_{b}^{\frac{1}{2}},$ (93)
$t_{w,i}$ $=[I_{p}-DR_{q}^{-1}H^{*}(\mu_{i})]^{-1}R_{c}^{\frac{1}{2}}.$ (94)
For DD-SW-BT:
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\sigma_{1})I_{m},\cdots,2\text{Real}(\sigma_{v})I_{m}\big),$ (95)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\mu_{1})I_{p},\cdots,2\text{Real}(\mu_{w})I_{p}\big),$ (96)
$T_{w}$ $\approx\text{blkdiag}\big([G(\mu_{1})]^{-*},\cdots,[G(\mu_{w})]^{-*}\big).$ (97)
For DD-BST:
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\sigma_{1})I_{m},\cdots,2\text{Real}(\sigma_{v})I_{m}\big),$ (98)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\mu_{1})[I_{p}-t_{w,1}^{*}\mathscr{H}(\mu_{1})R_{\text{X}}^{-1}\mathscr{H}^{*}(\mu_{1})t_{w,1}]^{-1},\cdots,$
$\hskip 56.9055pt2\text{Real}(\mu_{w})[I_{p}-t_{w,w}^{*}\mathscr{H}(\mu_{w})R_{\text{X}}^{-1}\mathscr{H}^{*}(\mu_{w})t_{w,w}]^{-1}\big),$ (99)
$T_{w}$ $\approx\text{blkdiag}\big(t_{w,1},\cdots,t_{v,w}\big),$ (100)
where
$\mathscr{H}(\mu_{i})$ $=\begin{cases}\frac{2\text{Real}(\sigma_{i})}{\mu_{i}-\sigma_{i}}\big[H(\sigma_{i})-H(\mu_{i})\big]H^{*}(\sigma_{i})+H(\mu_{i})D^{T}&\text{if }\sigma_{i}\neq\mu_{i}\\ 2\text{Real}(\sigma_{i})H^{\prime}(\sigma_{i})H^{*}(\sigma_{i})+H(\mu_{i})D^{T}&\text{if }\sigma_{i}=\mu_{i}\end{cases},$ (101)
$t_{w,i}$ $=\Big[R_{\text{X}}+\mathscr{H}^{*}(\mu_{i})\Big]^{-1}R_{\text{X}}^{\frac{1}{2}}.$ (102)
Using these formulas, the algorithms proposed in this paper can be efficiently implemented. Even if samples of $G(s)$ are available along the $j\omega$-axis, one can set a near-zero damping value and still use the above formulas and algorithms without significant loss of accuracy.
5 Numerical Example
This section evaluates the numerical performance of the proposed algorithms using the same $400^{th}$-order RLC circuit model considered in [3]. The MATLAB codes to reproduce the results are provided in [41]. All simulations were performed in MATLAB R2021b on a laptop with a 2 GHz Intel i7 processor and 16 GB of RAM.
Interpolation points for transfer function sampling are computed using formula (78). The imaginary parts $\omega_{i,\sigma}$ of the interpolation points $\sigma_{i}$ consist of 50 logarithmically spaced frequencies between $10^{-1}$ and $10^{4}$. The real parts are determined by setting $\zeta_{i,\sigma}=10^{-5}$. The points $\mu_{i}$ are identical to $\sigma_{i}$. Transfer function samples are obtained numerically from the state-space realization of the RLC model provided in [3]. The quantities $\sqrt{\lambda_{i}(P_{\text{LQG}}Q_{\text{LQG}})}$, $\sqrt{\lambda_{i}(P_{\mathcal{H}_{\infty}}Q_{\mathcal{H}_{\infty}})}$, $\sqrt{\lambda_{i}(P_{\text{PR}}Q_{\text{PR}})}$, $\sqrt{\lambda_{i}(P_{\text{BR}}Q_{\text{BR}})}$, $\sqrt{\lambda_{i}(PQ_{\text{SW}})}$, and $\sqrt{\lambda_{i}(PQ_{\text{S}})}$ are referred to as Hankel-like singular values.
Figures 16 compare the Hankel-like singular values and the relative error $\frac{\|G(s)-\hat{G}(s)\|_{\mathcal{H}_{\infty}}}{\|G(s)\|_{\mathcal{H}_{\infty}}}$. It can be seen that the $25^{th}$-order ROMs generated by intrusive methods (LQG-BT, $\mathcal{H}_{\infty}$-BT, PR-BT, BR-BT, SW-BT, and BST) and their non-intrusive counterparts (DD-LQG-BT, $D$-$\mathcal{H}_{\infty}$-BT, DD-PR-BT, DD-BR-BT, DD-SW-BT, and DD-BST) accurately capture the $20$ most dominant Hankel-like singular values. Moreover, the proposed data-driven algorithms achieve accuracy comparable to that of the intrusive methods for ROMs of orders ranging from $1$ to $25$.
(a) Hankel-like Singular Values Comparison
(b) Relative Error Comparison
Figure 1: Performance Comparison between LQG-BT and $D$-LQG-BT
(a) Hankel-like Singular Values Comparison
(b) Relative Error Comparison
Figure 2: Performance Comparison between $\mathcal{H}_{\infty}$-BT and $D$-$\mathcal{H}_{\infty}$-BT
(a) Hankel-like Singular Values Comparison
(b) Relative Error Comparison
Figure 3: Performance Comparison between PR-BT and $D$-PR-BT
(a) Hankel-like Singular Values Comparison
(b) Relative Error Comparison
Figure 4: Performance Comparison between BR-BT and $D$-BR-BT
(a) Hankel-like Singular Values Comparison
(b) Relative Error Comparison
Figure 5: Performance Comparison between SW-BT and $D$-SW-BT
(a) Hankel-like Singular Values Comparison
(b) Relative Error Comparison
Figure 6: Performance Comparison between BST and $D$-BST
6 Conclusion
This paper presents an interpolation-based implementation of the RADI algorithm for approximating large-scale Riccati equations. Building on this foundation, we propose several data-driven interpolation algorithms that preserve critical system properties, including positive-realness, bounded-realness, and minimum-phase characteristics. Furthermore, we develop data-driven implementations of LQG-BT and $\mathcal{H}_{\infty}$-BT, enabling the design of LQG and $\mathcal{H}_{\infty}$ controllers for plants without requiring state-space realizations. We also introduce data-driven versions of PR-BT, BR-BT, SW-BT, and BST algorithms. The paper addresses potential numerical implementation challenges and provides effective solutions. Numerical experiments demonstrate that the proposed algorithms achieve performance comparable to their intrusive counterparts.
Acknowledgment
We are deeply grateful to Ion Victor Gosea at the Max Planck Institute for Dynamics of Complex Technical Systems in Magdeburg, Germany, for his patient responses to our numerous questions about the Loewner framework and for his valuable feedback. We are also thankful to Prof. Patrick Kürschner at Leipzig University of Applied Sciences for his patient responses to our numerous questions about ADI method. This work is supported by the National Natural Science Foundation of China under Grants No. 62350410484 and 62273059, and in part by the High-end Foreign Expert Program No. G2023027005L granted by the State Administration of Foreign Experts Affairs (SAFEA).